### v2

# Composition of Scalar Mesons and Their Effects on Nuclear Matter Properties in an Extended Linear Sigma Model

Abstract
It has been argued that the iso-scalar and iso-vector mesons play significant roles in nuclear matter and neutron star structures. We improve the extended linear sigma model with baryons, proposed in our previous work, by introducing the flavor structures constructed from antisymmetric tensors of chiral representations to study these physics. The parameter space of this model is refined with well-reproduced nuclear matter properties at saturation density by the lowest order Lagrangian, ensuring consistency with vacuum results, such as $f_{\pi}\approx 134\,\text{MeV}$. The anticipated plateau-like behaviors of the symmetry energy are predicted at intermediate densities, which is crucial for the consistency of GW170817 and the neutron skin thickness of $\text{Pb}^{208}$. Subsequently, neutron star structures are calculated using several parameter sets, and the results for the nuclear matter properties at saturation density align with empirical values. It is found that the neutron star structures are sensitive to the couplings between the iso-vector $a_{0}$ meson and nucleons and the four-vector meson couplings: small values of both are favorable. Meanwhile, nuclear matter properties at saturation density favor larger values of the latter and are not sensitive to the former. This signifies the statistical significance of neutron star observations when obtaining realistic chiral effective field theories or models at various densities. The parameter set favored by neutron star observations also aligns the behavior of the sound velocity with the conformal limit at high densities relevant to cores of massive stars. It is hoped that the results of this work can guide future studies on the relationship between the microscopic symmetry of strong interactions and macroscopic phenomena.
I Introduction
Nuclear matter (NM), a system dominated by the nonperturbative quantum chromodynamics (QCD), has been studied for several decades. But there are still great uncertainties about its properties, especially at compact star densities, due to our little knowledge on the phase structure of strong interaction matter Fukushima and Hatsuda (2011); Lattimer and Prakash (2016); Chen (2017); Baym et al. (2018); Gil et al. (2019); Ma and Rho (2020); Li et al. (2019); Sorensen et al. (2024). The properties of NM are highly sensitive to details of nucleon interactions. Therefore, a reliable model of NM is not only significant for revealing neutron star (NS) structures Lattimer and Prakash (2001); Özel and Freire (2016), but also can shed a light on characters of the hadrons involved.
In modeling NM, based on the phenomenological one-boson exchange for nucleon interactions Holinde and Machleidt (1975); Erkelenz (1974); Nagels et al. (1978); Machleidt (1989), it has long been recognized that the scalar meson resonances, such as $\sigma,f_{0}$ and $a_{0}(\delta)$, play indispensable roles by using the mean field approach (MFA) Walecka (1974); Serot and Walecka (1986, 1997) [^1] Hereafter, following the Particle Data Group, we denote the iso-vector scalar mesons as $a_{0}$. Without the iso-scalar scalar mesons $\sigma/f_{0}$, which provide the attractive force between nucleons, it’s hard to obtain the empirical values of the saturation density $n_{0}\simeq 0.16$ fm-3 and binding energy $E_{0}\simeq 16.0$ MeV Holinde and Machleidt (1975); Erkelenz (1974); Nagels et al. (1978); Machleidt (1989). And the self-interactions of $\sigma$ meson is necessary to reduce the nuclear incompressibility coefficient of symmetric nuclear matter Boguta and Bodmer (1977) and other empirical values Motohiro et al. (2015). In addition to the iso-scalar scalar mesons, the iso-vector scalar mesons are important for describing the asymmetric nuclear matter and hence the properties of neutron stars Kubis and Kutschera (1997); Hofmann et al. (2001); Liu et al. (2002); Chen et al. (2007); Roca-Maza et al. (2011); Wang et al. (2014); Typel and Alvear Terrero (2020); Kong et al. (2025). Furthermore, it is found that the scalar-iso-scalar scalar meson couplings, e.g., $\sigma a_{0}^{2}$ and $\sigma^{2}a_{0}^{2}$ have large effects on both the symmetry energy $E_{\rm sym}$ and its slope $L$ Zabari et al. (2019); Kubis et al. (2020); Miyatsu et al. (2022); Li et al. (2022). Explicitly, they soften $E_{\rm sym}$ at intermediate densities but stiffen $E_{\rm sym}$ at high densities.
Normally, in the models of NM, the scalar mesons are included with respect to that quantum numbers, such as parity, spin and isospin, less attention has been paid to other properties related to QCD. From QCD point of view, the structures of scalar mesons are still under debate, especially the light ones (see, e.g., Close and Tornqvist (2002) for a review). From the mass ordering of the scalar mesons below $1$ GeV, it is difficult to regard them as the pure two-quark states but an admixture of the two- and four-quark configurations. Considering that the two quark and four quark components of scalar mesons have different axial transformation properties, we recently proposed an extended linear sigma model (ELSM) with baryons, taken as diquark-quark configurations, (bELSM) to study the properties of NM Ma and Ma (2024).
In bELSM, the masses of the mesons and baryons are generated from the spontaneous chiral symmetry breaking. Due to the conservation of axial transformation, only the 2-quark component of the scalar mesons couples to the baryons at the lowest order, which may provide deeper insights into the constituents of the scalar mesons from the NM and NS properties. The $\rm U(1)_{A}$ anomaly is involved in the ’t Hooft determinant Rosenzweig et al. (1980); Kawarabayashi and Ohta (1980); Hsu et al. (1998); Fariborz et al. (2005). Another important feature of the bELSM for the phenomenological analysis is the inclusion of all possible types of mesons below $1\ \rm GeV$ and hyperons which significantly affect NS structures Gal et al. (2016).
After pinning down the NM properties at saturation density $n_{0}$, a plateau-like structure of the symmetry energy, $E_{\rm sym}$, at intermediate densities, $\sim 2n_{0}$, was found in the bELSM. This is consistent with the analysis based on various of approaches in Refs. Paeng et al. (2017); Ma and Rho (2019); Ma et al. (2019); Ma and Rho (2021); Lee et al. (2022); Zabari et al. (2019); Miyatsu et al. (2022); Li et al. (2022), and satisfies the constraints of $\rm{}^{208}Pb$ Adhikari et al. (2021); Reed et al. (2021) and tidal deformations of NS from GW170817 Abbott et al. (2016). We also investigated the effects of parameter variations on different coefficients of Taylor expansions of energy per-nucleon, $E=\mathcal{E}/n$, by density, $n=n_{n}+n_{p}$, and asymmetry, $\alpha=(n_{n}-n_{p})/n$, to provide further guidance on the interaction parameterization of NM:
$E(n,\alpha)$ $\simeq$ $E(n_{0},0)+E_{\rm sym}(n_{0})\alpha^{2}+L(n_{0})\alpha^{2}\chi+\frac{K(n_{0})}{2!}\chi^{2}+\frac{J(n_{0})}{3!}\chi^{3}+\mathcal{O}\left(\alpha^{4}\right)+\mathcal{O}\left(\chi^{4}\right)+\mathcal{O}\left(\alpha^{2}\chi^{2}\right)$ (1)
$\simeq$ $E_{0}+\left.\frac{\alpha^{2}}{2}\frac{\partial^{2}E(n_{0},\alpha)}{\partial\alpha^{2}}\right|_{\alpha=0}+\alpha^{2}\chi\left.\left(3n\frac{\partial E_{\mathrm{sym}}(n)}{\partial n}\right)\right|_{n=n_{0}}+\frac{\chi^{2}}{2!}\left.\left(9n^{2}\frac{\partial^{2}E_{0}(n)}{\partial n^{2}}\right)\right|_{n=n_{0}}$
${}+\frac{\chi^{3}}{3!}\left.\left(27n^{3}\frac{\partial^{3}E_{0}(n)}{\partial n^{3}}\right)\right|_{n=n_{0}}+\mathcal{O}\left(\alpha^{4}\right)+\mathcal{O}\left(\chi^{4}\right)+\mathcal{O}\left(\alpha^{2}\chi^{2}\right)\ ,$
where $\chi\equiv\left(n-n_{0}\right)/3n_{0}$. $L(n_{0})$, $K(n_{0})$, and $J(n_{0})$ refer to the symmetry energy slope, incompressibility, and skewness of NM at $n_{0}$, respectively.
However, when calculating the $E_{\rm sym}(n)$ using bELSM, the plateau-like structure is found to be lower than the results in Ref. Miyatsu et al. (2022); Li et al. (2022), leading to a loose structure of NS with tidal deformation $\Lambda_{1.4}>1000$. This is due to the couplings related to $\rho$ and $\omega$ mesons being constrained by the $\rm U(3)_{V}$ symmetry, which is not entirely realistic, especially in nuclear medium Paeng et al. (2012). Therefore, improving the bELSM to resolve these problems and achieve a better description of the NM properties and NS structures is the purpose of this work.
When constructing the interactions, especially in the baryonic sector, only operators like ${\rm Tr}(\bar{B}\gamma_{\mu}V^{\mu}B)$ were considered in Ref. Ma and Ma (2024), which make the couplings $g_{\rho NN}$ and $g_{\omega NN}$ the same. However, another type of operator, $\epsilon_{abc}\epsilon^{def}\bar{B}_{ad}\gamma_{\mu}B_{be}V^{\mu}_{cf}$ Papazoglou et al. (1998), which can split the couplings of the singlet and octet vector mesons to baryons. In this work, we extend the bELSM by including this kind of operators in both the mesonic and baryonic sectors at the lowest order and improve the parameter space of the original bELSM. The 12-quark terms, which were necessary to maintain reasonable behavior at high densities in the work of Ref. Ma and Ma (2024), can be removed, enhancing the rationality of the power-counting in the bELSM. The vacuum expectation values (VEVs) of the scalar meson fields are close to the results of Ref. Fariborz et al. (2008), leading to a reasonable $f_{\pi}$ (quark condensate) and low-energy pion scattering results.
With the improved bELSM at the lowest order, the NM properties at $n_{0}$ and the bare masses of the mesons and baryons can be well reproduced. The bELSM is then applied to calculating NS structures. It is found that NS structures are sensitive to the NM properties at $n_{0}$, especially the incompressibility, $K(n_{0})$, which also affects the sound velocity (SV) behavior at various densities. The relationship between NS and SV is very subtle, as detailed in Refs. Bedaque and Steiner (2015); Zhang et al. (2024a). When $K(n_{0})$ falls within empirical constraints, the SV softens significantly, causing the maximum mass of NS to be lower than the observed value, $2.14\pm 0.10M_{\odot}$ Cromartie et al. (2019). To meet the constraints of the maximum mass of NS and the mass-radius (M-R) relation of NS from GW170817 Abbott et al. (2016), $K(n_{0})$ should be larger than the empirical constraints, approximately $500\ \rm MeV$. This $K(n_{0})$ also results in an interesting behavior of SV, approaching the conformal limit at high densities, $n\sim 5n_{0}$, whereas other SVs from smaller $K(n_{0})$ will be too soft. This highlights the statistical significance of NS observations when obtaining realistic chiral effective field theories ($\chi$EFTs) or models Guo et al. (2024a). In addition, the influence of portions in the 2-quark and 4-quark constituents of the scalar mesons on NS structures and NM properties is also shown and their physical magnitudes are estimated. These discussions setup a relationship between microscopic symmetry and macroscopic phenomena.
The rest of this article is organized as follows: In Sec. II, we first introduce the improved bELSM at the lowest order under the relativistic mean field (RMF) approximation. Then we study the properties of the scalar mesons on the NM properties and NS structures in detail. In Sec. III, the discussion on the results of the current work and the outlook are provided. The details of the model construction and approximation are given in Appendix.
II Extended linear sigma model at densities
II.1 Theoretical framework
In this work, we consider the lowest-order bELSM given in Ref. Ma and Ma (2024) including terms up to 8-quark lines with a single trace and operators of dimension-4. In addition, as mentioned in the Introduction, we consider the terms with three-rank Levi-Civita tensor in flavor space. The Lagrangian contributing under RMF approximation is given by
$\mathcal{L}=\mathcal{L}_{\rm B}+\mathcal{L}_{\rm M}+\mathcal{L}_{\rm V}$ (2)
where
$\mathcal{L}_{\rm B}$ $=$ ${\rm Tr}\left(\bar{B}i\not{\partial}B\right)+c{\rm Tr}\left(\bar{B}V\gamma^{0}B\right)$ (3a)
${}-g{\rm Tr}\left(\bar{B}S^{\prime}B\right)+h\epsilon_{abc}\epsilon^{def}\bar{B}_{ad}\gamma^{0}B_{be}V_{cf}$
${}-e\epsilon_{abc}\epsilon_{def}\bar{B}_{ad}B_{be}S_{cf}^{\prime}\ ,$
$\mathcal{L}_{\rm M}$ $=$ $c_{2}{\rm Tr}S^{\prime 2}-d_{2}{\rm Tr}\hat{S}^{\prime 2}-c_{4}{\rm Tr}S^{\prime 4}$ (3b)
${}-2e_{3}\epsilon_{abc}\epsilon_{def}S_{ad}^{\prime}S_{be}^{\prime}\hat{S}_{cf}^{\prime}\ ,$
$\mathcal{L}_{\rm V}$ $=$ $\tilde{h}_{2}{\rm Tr}\left(S^{\prime 2}V^{2}\right)+\tilde{g}_{3}{\rm Tr}V^{4}$ (3c)
${}+a_{1}\epsilon_{abc}\epsilon_{def}V_{ad}V_{be}\left(S^{\prime 2}\right)_{cf}\ .$
The $h,\ e$ and $a_{1}$ terms are the terms which were not considered in Ref. Ma and Ma (2024). The details of the model construction are given in App. A.
In Lagrangian (2), $S^{\prime}$ is the two-quark configuration of scalar states which can be decomposed as
$S^{\prime}$ $=$ $\frac{1}{\sqrt{2}}(\lambda_{8}f_{0}^{\prime}+\lambda_{3}a_{0}^{\prime})+\frac{1}{\sqrt{3}}I\sigma^{\prime},$ (4)
with $\lambda_{3}$ and $\lambda_{8}$ being the 3rd and 8th components of Gell-Mann matrices, and $I$ as the identity matrix. The similar decomposition applies to the four-quark configuration $\hat{S}^{\prime}$. $V$ is the zero component of vector fields which survives in the MFA
$V$ $=$ $\frac{1}{2}\left(\lambda_{3}\rho+\frac{1}{\sqrt{3}}\lambda_{8}\omega\right)+\frac{1}{3}I\omega.$ (5)
When restricting to the two-flavor case, the baryon octet is reduced to
$B\rightarrow\left(\begin{array}[]{ccc}0&0&p\\ 0&0&n\\ 0&0&0\end{array}\right)\ .$ (6)
So, with $p$ and $n$, there are iso-scalar scalar configurations $\sigma^{\prime}/\hat{\sigma}^{\prime}$, $f_{0}^{\prime}/\hat{f}_{0}^{\prime}$ and iso-vector scalar configurations $a_{0}^{\prime}/\hat{a}_{0}^{\prime}$ in model (2). After chiral symmetry breaking, the mixing of these scalar configurations yields the physical states of scalar mesons. The spontaneous chiral symmetry breaking is achieved by the vacuum expectation values of both 2-quark and 4-quark fields: $\langle\sigma^{\prime}\rangle=\sqrt{3}\alpha$ and $\langle\hat{\sigma}^{\prime}\rangle=\sqrt{3}\beta$, which are determined by the minimum of the potential $\mathcal{L}_{\rm M}$:
$\left\langle\frac{\partial V_{\rm M}}{\partial\sigma^{\prime}}\right\rangle$ $=$ $2\alpha\left(-c_{2}+2c_{4}\alpha^{2}+4e_{3}\beta\right)=0\ ,$
$\left\langle\frac{\partial V_{\rm M}}{\partial\hat{\sigma}^{\prime}}\right\rangle$ $=$ $2\left(d_{2}\beta+2e_{3}\alpha^{2}\right)=0\ ,$ (7)
where the three-flavor symmetry is used. The physical states $\sigma$, $a_{0}$, and $f_{0}$ are obtained by the mixing of the 2-quark and 4-quark configurations, via
$\sigma=\cos\theta_{0}\sigma^{\prime}+\sin\theta_{0}\hat{\sigma}^{\prime}\ ,$
$a_{0}=\cos\theta_{8}a_{0}^{\prime}+\sin\theta_{8}\hat{a_{0}}^{\prime}\ ,$
$f_{0}=\cos\theta_{8}f_{0}^{\prime}+\sin\theta_{8}\hat{f_{0}}^{\prime}\ ,$ (8)
which are determined by the diagonalization of the mass matrix of the scalar mesons. As one can see, there are two sets of physical scalar mesons: one contains the lowest-lying scalar mesons $\sigma$, $a_{0}$, and $f_{0}$, while the other set refers to the excitation states in the corresponding channels, whose contribution to the equation of state (EoS) of NM is negligible but their spectra are considered to settle down the free parameters.
II.2 Nuclear matter properties and neutron star structures
Using the relations in Eq. (7), the degrees of freedom of the parameter space are reduced to $11$, explicitly, $\alpha$, $\beta$, $c_{4}$, $e_{3}$, $\tilde{h}_{2}$, $\tilde{g}_{3}$, $a_{1}$, $c$, $g$, $h$, and $e$. The physical quantities that can be used to constrain these parameters are the bare masses of the scalar mesons and nucleons, as well as the NM properties around saturation density, $n_{0}$. The empirical constraints and corresponding fitted values are shown in Tabs. 2 and 3, and the fitted parameters are listed in Tab. 1.
Table 1: Values of the parameters used in this work. $\alpha$ and $\beta$ are in units of MeV, and $e_{3}$ is in units of GeV. The other parameters are dimensionless. The parameter sets are denoted as el-g30g, el-g30e, el-g30eg, el-g350eg, el-g3100eg, and el-g3150eg, respectively. ’el’ refers to the bELSM; ’g3’ refers to the four-vector meson coupling in Eq. (3c), with the number following it denoting the corresponding magnitude; ’e’ or ’g’ indicates that the coupling between baryons and scalar mesons is mainly from ’e’ or ’g’ terms in Eq. (3a), and ’eg’ indicates that the contributions of the ’e’ and ’g’ terms are similar.
$\alpha$ $\beta$ $e_{3}$ $c_{4}$ $\tilde{h}_{2}$ $\tilde{g}_{3}$ $c$ $g$ $a_{1}$ $h$ $e$
el-g30g $61.1$ $24.7$ $-2.06$ $44.0$ $80.1$ $1.59$ $-0.792$ $15.4$ $4.25$ $11.4$ $-0.0270$
el-g30e $61.1$ $24.4$ $-2.05$ $43.6$ $80.0$ $0.542$ $11.4$ $0.234$ $4.17$ $-0.790$ $15.1$
el-g30eg $61.4$ $26.4$ $-2.10$ $45.6$ $79.3$ $0.397$ $9.51$ $6.54$ $4.10$ $-2.61$ $8.75$
el-g350eg $61.2$ $25.6$ $-2.10$ $44.4$ $79.9$ $51.5$ $10.1$ $6.35$ $4.14$ $-2.65$ $9.00$
el-g3100eg $60.8$ $24.0$ $-2.09$ $42.4$ $80.8$ $100$ $10.6$ $7.10$ $4.19$ $-2.88$ $8.34$
el-g3150eg $60.7$ $24.3$ $-2.11$ $42.0$ $81.0$ $150$ $11.1$ $7.15$ $4.18$ $-3.05$ $8.30$
Table 2: The spectra of hadrons are in units of $\mathrm{MeV}$. The empirical values are chosen as the real parts of the corresponding resonance T-matrix poles or masses from Ref. Navas et al. (2024). The excited state $f_{0}^{\prime}$ is chosen to be $f_{0}(1370)$ with $a_{0}^{\prime}$ being $a_{0}(1450)$ and $\sigma^{\prime}$ being $f_{0}(1500)$. The constraints of $f_{0}(1370)$ and $a_{0}(1450)$ are combined together, due to $\rm U(3)_{V}$ symmetry. The definition of the parameter set is shown in Tab. 1.
$m_{N}$ $m_{\sigma}$ $m_{f_{0}(a_{0})}$ $m_{f_{0}^{\prime}(a_{0}^{\prime})}$ $m_{\sigma^{\prime}}$ $m_{\rho}$ $m_{\rho}$
Empirical $938$-$940$ $400$-$800$ $960$-$1010$ $1250$-$1500$ $1430$-$1530$ $761$-$765$ $782$-$783$
el-g30g $939$ $522$ $989$ $1480$ $1510$ $763$ $783$
el-g30e $939$ $525$ $991$ $1480$ $1510$ $763$ $783$
el-g30eg $939$ $498$ $983$ $1500$ $1520$ $763$ $783$
el-g350eg $939$ $485$ $991$ $1470$ $1510$ $763$ $783$
el-g3100eg $939$ $502$ $994$ $1470$ $1510$ $763$ $783$
el-g3150eg $939$ $485$ $991$ $1470$ $1510$ $763$ $783$
Table 3: Nuclear matter properties at saturation density $n_{0}$. The empirical values and their error bars are set according to analyses from Refs. Sedrakian et al. (2023); Piekarewicz (2004); Colo et al. (2004); Khan and Margueron (2012); Dutra et al. (2012, 2014); Oertel et al. (2017); Tews et al. (2017); Wang et al. (2018); Xie and Li (2019); Choi et al. (2021); Grams et al. (2022). The $n_{0}$ is in units of $\rm fm^{-3}$ and the other properties are in units of $\rm MeV$. The definition of the parameter set is shown in Tab. 1.
$n_{0}$ $E_{0}$ $E_{\mathrm{sym}}(n_{0})$ $J(n_{0})$ $L(n_{0})$ $K(n_{0})$
Empirical $0.155\pm 0.050$ $-16.0\pm 1.0$ $31.7\pm 3.2$ $-200\pm 600$ $58.7\pm 28.1$ $250\pm 50$
el-g30g $0.155$ $-14.6$ $30.9$ $451$ $83.6$ $418$
el-g30e $0.155$ $-14.6$ $31.6$ $479$ $85.8$ $419$
el-g30eg $0.155$ $-14.6$ $30.1$ $421$ $82.2$ $415$
el-g350eg $0.155$ $-15.2$ $30.9$ $-392$ $80.7$ $370$
el-g3100eg $0.155$ $-15.4$ $31.4$ $-1020$ $71.7$ $317$
el-g3150eg $0.155$ $-15.6$ $31.6$ $-1470$ $63.7$ $253$
It can be seen from Tab. 2 that all the parameter sets reproduce hadron spectra consistent with the results from the Particle Data Group Navas et al. (2024). Regarding the NM properties at $n_{0}$ shown in Tab. 3, the saturation density is consistently fixed at $0.155~{}\rm fm^{-3}$, and $E_{\rm sym}$ aligns with the empirical value of $(30.9\pm 1.9)~{}\rm MeV$. However, the other properties decrease with the magnitude of $\tilde{g}_{3}$. It appears that $\tilde{g}_{3}$ should be approximately $150$ to satisfy the constraints of NM properties.
When we extend our study beyond the saturation density to a broader density range to examine NS structures, these parameter sets result in significant differences. The M-R relations of NSs are shown in Fig. 1, which is computed using the Tolman-Oppenheimer-Volkoff (TOV) equation Tolman (1939); Oppenheimer and Volkoff (1939) with pure hadron EOSs from different models.
(a) Comparison among TM1 Sugahara and Toki (1994), FSU-$\delta 6.7$ Li et al. (2022), and el-g30eg.
(b) Comparison among cases with different magnitudes of $e$ and $g$ flavor structures.
(c) Compasrion among cases with different magnitude of $\tilde{g}_{3}$ terms.
Figure 1: The M-R relation of NSs. The constraints MSP J0740+6620 are from Ref. Cromartie et al. (2019) and GW170817 are from Ref. Abbott et al. (2018). Both are at $95\%$ confidence level.
As we can see in Fig. 1(a), the model without the $a_{0}$ meson cannot reach the GW170817 constraint due to its overly loose structure of NSs. The models with the $a_{0}$ meson, both el-g30eg and FSU-$\delta 6.7$, can satisfy the GW170817 and MSP J0740+6620 constraints simultaneously, highlighting the importance of the $a_{0}$ meson—an iso-vector scalar meson in NS studies. From Fig. 1(b), it is evident that only EOSs with similar contributions from the $e$ and $g$ terms can meet the observational constraint of MSP J0740+6620; otherwise, the $M_{\rm max}$ will be slightly lower. However, the properties of NM at $n_{0}$ are not sensitive to the contributions of the $e$ and $g$ terms. Regarding the $\tilde{g}_{3}$ terms, as shown in Fig. 1(c), the $M_{\rm max}$ of NSs is sensitive to the magnitude of $\tilde{g}_{3}$: a larger $\tilde{g}_{3}$ coupling—four vector meson couplings—leads to a smaller $M_{\rm max}$ of NSs, even though the NM properties at $n_{0}$ are more consistent with empirical values.
To further understand the effects of the meson-nucleon coupling on the NS structures, we combine the parameters in various of model to the meson-baryon couplings defined by
$\mathcal{L}_{\rm RMF}^{\rm OBE}$ $=$ $g_{\sigma NN}\bar{N}\sigma N+g_{aNN}\bar{N}a_{0}\tau_{3}N$ (9)
${}+g_{fNN}\bar{N}f_{0}N+g_{\omega NN}\bar{N}\omega\gamma_{0}N$
${}+g_{\rho NN}\bar{N}\rho\tau_{3}\gamma_{0}N,$
where $N=\binom{p}{n}$ and $\tau_{3}$ is the 3rd component of Pauli matrix. The values of the couplings are list in Table 4.
Table 4: The Values of meson-baryon coupling in various of models.
$g_{\sigma NN}$ $g_{aNN}$ $g_{fNN}$ $g_{\omega NN}$ $g_{\rho NN}$
TM1 Sugahara and Toki (1994) $-10.0$ $-12.6$ $-4.63$
FSU-$\delta$6.7 Li et al. (2022) $10.2$ $6.7$ $-13.4$ $-7.27$
el-g30g $-6.17$ $5.13$ $2.95$ $-6.09$ $5.30$
el-g30e $-6.20$ $-5.03$ $3.00$ $6.09$ $5.30$
el-g30eg $-5.97$ $-0.671$ $2.68$ $6.06$ $3.45$
el-g350eg $-6.12$ $-0.852$ $2.85$ $6.37$ $3.71$
el-g3100eg $-6.36$ $-0.442$ $3.19$ $6.73$ $3.85$
el-g3150eg $-6.38$ $-0.413$ $3.20$ $7.09$ $4.04$
In our bELSM, the magnitudes of the $e$ and $g$ terms determine $g_{aNN}$, while $g_{\sigma NN}$ and $g_{fNN}$ remain almost the same across different parameter sets. However, the M-R relation of NSs is sensitive to $g_{aNN}$: a larger $g_{aNN}$ leads to a smaller $M_{\rm max}$ for NSs, as shown in Fig. 1(b). This is not the case in FSU-$\delta 6.7$, which does not consider $f_{0}$. In bELSM, all possible terms involving the $\delta$ meson at the lowest order are considered, including all three- and four-point interactions: self-couplings, couplings to vector mesons, and couplings to scalar mesons. These terms become increasingly important with density, especially those relate to vector mesons. By including all possible multi-meson interactions at the lowest order, respecting the chiral symmetry pattern at low energies, the $g_{\sigma NN}$ and $g_{\omega NN}$ at densities are slightly smaller than those in Walecka-type models Holinde and Machleidt (1975); Erkelenz (1974); Nagels et al. (1978); Machleidt (1989); Sugahara and Toki (1994); Li et al. (2022), with $g_{\rho NN}$ remaining almost the same.
Furthermore, the $E_{\rm sym}(n)$ across the relevant densities is shown in Fig. 2.
(a) Comparison among TM1, FSU-$\delta 6.7$, and el-g30eg. TM1 is from Ref. Sugahara and Toki (1994). FSU-$\delta 6.7$ is from Ref. Li et al. (2022).
(b) Comparison among cases with different magnitudes of $e$ and $g$ flavor structures.
(c) Comparison among cases with different magnitudes of $\tilde{g}_{3}$ terms.
Figure 2: The $E_{\rm sym}(n)$ for different cases.
As shown in Fig. 2(a), the $E_{\rm sym}(n)$ from el-g30eg and FSU-$\delta$6.7 is soft at intermediate densities, around $2n_{0}$, while stiff at subsaturation densities, around $2/3n_{0}$, to meet the constraints of GW170817 Abbott et al. (2018) and the neutron skin thickness of ${}^{208}\rm Pb$ Adhikari et al. (2021); Reed et al. (2021). However, the $E_{\rm sym}(n)$ from TM1 is almost linearly dependent on the density, which is too stiff at intermediate densities, leading to a too loose NS compared to the observations of GW170817, as shown in Fig. 1(a). This highlights the necessity of including the $a_{0}$ meson in the NS EOS, resulting in a plateau-like structure of $E_{\rm sym}(n)$ at intermediate densities. However, a large $g_{aNN}$ will lead to a decreasing $E_{\rm sym}(n)$ at high densities, as shown in Fig. 2(b), which results in a decrease in the $M_{\rm max}$ of NSs, as mentioned in Fig. 1(b). When comparing the $E_{\rm sym}(n)$ from different $\tilde{g}_{3}$ terms in Fig. 2(c), it seems that $E_{\rm sym}(n)$ is not significantly suppressed by the $\tilde{g}_{3}$ term, but the $M_{\rm max}$ of NSs is even smaller than in the case with a large $g_{aNN}$. The reason lies in the SV behavior at high densities, as presented in Fig. 3.
(a) Comparison among TM1, FSU-$\delta 6.7$, and el-g30eg. TM1 is from Ref. Sugahara and Toki (1994). FSU-$\delta 6.7$ is from Ref. Li et al. (2022).
(b) Comparison among cases with different magnitudes of $e$ and $g$ flavor structures.
(c) Comparison among cases with different magnitudes of $\tilde{g}_{3}$ terms.
Figure 3: The $v_{s}^{2}$ from different cases. The conformal limit is due to the asymptotic freedom of QCD at high energies.
It can be observed that a large $\tilde{g}_{3}$ makes $v_{s}^{2}$ softer compared to $g_{aNN}$ suppression, as seen in Fig. 3(b) and Fig. 3(c). The soft SV behavior, especially when not approaching the conformal limit throughout the densities, predicts a small $M_{\rm max}$ for NSs, which cannot meet the constraints of MSP J0740+6620.
Now, let’s examine the results in Fig. 1(a):
  • The SV of TM1 is the stiffest throughout the densities, but the predicted NS is the loosest.
  • The SV of FSU-$\delta 6.7$ is the softest at low densities but becomes stiff at intermediate/high densities, exceeding the conformal limit $v_{s}^{2}=1/3$, leading to the most compact NS.
  • The SV of el-g30eg is similar to TM1 at low densities but approaches the conformal limit at high densities, with NS results consistent with the observations of both GW170817 and MSP J0740+6620.
All these indicate that the SV behavior should be stiff enough at high densities to make the $M_{\rm max}$ of NSs reach at least the observed value of MSP J0740+6620, and the SV behavior should be soft enough at intermediate densities to make the NSs compact enough to meet the constraints of GW170817. This reflects the subtle relationship between NS structures and SV.
If we only consider the NM properties and the NS structures, both el-g30eg and FSU-$\delta 6.7$ can meet the phenomenological constraints, with FSU-$\delta 6.7$ appearing more favored. However, bELSM reflects the symmetry pattern of QCD at low energies, while the Walecka-type models lack such consideration but introduce possible couplings artificially. Moreover, considering the SV behavior at high densities, el-g30eg allows the SV to approach the conformal limit. This technically stems from including all possible terms about the related baryons and mesons at the lowest order, but the deeper reason may be its close relation to QCD. All the M-R relations in this work are calculated with parameters determined by the NM properties at $n_{0}$, but the parameter space and the resulting NS structures can be further improved via statistical analysis, including constraints from NS observations, as discussed in our past work Guo et al. (2024a).
II.3 Constituents of the scalar mesons
The above discussions suggest possible magnitudes of the 2-quark and 4-quark constituents of the scalar mesons. We list the percentage of the two-quark component $p_{2}=\left|\cos^{2}\theta\right|$ and four-quark component $p_{4}=\left|\sin^{2}\theta\right|$ defined by Eq. (8) in Tab. 5.
Table 5: The possible percentage of the 2-quark and 4-quark constituents of the scalar mesons.
$\sigma$ $\delta$
$p_{4}$ $p_{2}$ $p_{4}$ $p_{2}$
el-g30g $51.7\%$ $48.3\%$ $77.9\%$ $22.1\%$
el-g30e $51.2\%$ $48.7\%$ $77.2\%$ $22.8\%$
el-g30eg $54.2\%$ $45.8\%$ $81.5\%$ $18.5\%$
el-g350eg $52.4\%$ $47.6\%$ $77.3\%$ $20.7\%$
el-g3100eg $49.2\%$ $50.8\%$ $74.4\%$ $25.6\%$
el-g3150eg $48.9\%$ $51.1\%$ $74.3\%$ $25.7\%$
It is found that the $\sigma$ meson, a flavor singlet, is composed of 2-quark and 4-quark constituents in similar proportions, while the $\delta$ meson, a flavor octet, is primarily composed of 4-quark constituents. This conclusion is drawn after determining the bare masses of the relevant hadrons in Tab. 2 and the NM properties in Tab. 3. This finding is consistent with previous analyses in vacuum Fariborz et al. (2008). When increasing the magnitude of the $\tilde{g}_{3}$ term, the 4-quark constituent percentages of both the $\sigma$ and $\delta$ mesons decrease to find an optimal parameter set that meets empirical constraints. Conversely, when the contributions of the $e$ and $g$ terms are made similar, the 4-quark constituent percentages of both the $\sigma$ and $\delta$ mesons increase.
Regarding the parameter space found in our previous work Ma and Ma (2024), the current optimal parameter spaces are more favored, not only for their consistent NM properties at $n_{0}$ and reasonable EOS behaviors over a wider density region, as discussed above, but also for their prediction of the pion decay constant $f_{\pi}$. The pion decay constant can be obtained using the PCAC relation in the chiral limit (no explicit symmetry breaking due to quark mass) Fariborz et al. (2008): $f_{\pi}=2\sqrt{\alpha^{2}+\beta^{2}}$. The el-g30eg parameter set indicates that $f_{\pi}\approx 134~{}\rm MeV$, which is consistent with recent lattice QCD results Aoki et al. (2020), while $f_{\pi}\sim(60-70)~{}\rm MeV$ was obtained from the parameter space in our previous work. This again highlights the importance of the new flavor structures in the phenomenological analysis to achieve a realistic parameter space.
The $a_{0}$ meson plays a crucial role in NM studies, particularly for pure neutron matter at high densities Kubis and Kutschera (1997); Hofmann et al. (2001); Liu et al. (2002); Chen et al. (2010); Zabari et al. (2019); Li et al. (2022). The couplings related to the $a_{0}$ meson are traditionally introduced artificially in Walecka-type models and are usually fixed by the properties of NM around $n_{0}$. In this work, to reproduce the rational spectra of mesons and baryons, with respect to the chiral symmetry pattern of QCD at low energies, along with the NM properties at $n_{0}$, the couplings related to the $a_{0}$ meson are tightly constrained.
Effectively, we define the coupling through
$\mathcal{L}_{\rm RMF}$ $=$ $C_{\sigma\delta^{2}}\sigma a_{0}^{2}+C_{\sigma^{2}\delta^{2}}\sigma^{2}a_{0}^{2}.$ (10)
Then, according to Tab. 1 and Lagrangian (3b), the corresponding values of the couplings are listed in Tab. 6 for comparison.
Table 6: The values of $C_{\sigma\delta^{2}}$ and $C_{\sigma^{2}\delta^{2}}$ in different models.
$C_{\sigma\delta^{2}}$(MeV) $C_{\sigma^{2}\delta^{2}}$
Zabari-19 Zabari et al. (2019) ±$\pm 1.77$$\pm 0.004$
FSU-$\delta$6.7 Li et al. (2022) $2.63$
el-g30g $-1860$ $-9.40$
el-g30e $-1940$ $-9.71$
el-g30eg $-1480$ $-7.70$
el-g350eg $-1690$ $-8.75$
el-g3100eg $-2190$ $-11.0$
el-g3150eg $-2160$ $-11.0$
It can be seen that the magnitude of couplings between $a_{0}$ and $\sigma$ mesons are all large, differing significantly from the results of FSU-$\delta 6.7$ and Zabari-19, which are used for dense NM studies.
The larger magnitude of the couplings in Tab. 6 indicates that these multi-meson interactions contribute more to the EOS at both saturation density and high densities, and links the behavior of both regions more closely. This is the reason why we must increase the $K(n_{0})$ to achieve a reasonable NS structures, as shown Fig. 1(c). While the parameter spaces of FSU-$\delta 6.7$ and Zabari-19 make NM properties at $n_{0}$ do not change obviously after introducing these scalar couplings, they affect the EOS at high densities much. This leads to the fact that they can handle the NM properties at $n_{0}$ and NS structures separately, compared to the bELSM. However, since the scalar meson couplings in $\mathcal{L}_{\rm M}$ are determined by $\alpha$, $\beta$, $e_{3}$, and $c_{4}$, which are closely related to the quark constituents of scalar mesons listed in Tab. 5 and important physical values, such as $f_{\pi}$. In contrast, the Walecka-type models, as we are addressing, only introduce possible freedoms and choose operators artificially, which cannot be directly linked to QCD. Therefore, the results in Tab. 6 highlight the importance of considering the chiral symmetry pattern of QCD at low energies in phenomenological analysis and including operators systematically, especially for phenomenological studies at high densities.
III Conclusion and discussion
In this work, we continued to study NM properties and NS structures by extending the bELSM previousely proposed in Ref. Ma and Ma (2024) to include a new type of flavor structure which is constructed with an antisymmetric tensor of the corresponding baryon and meson representations. With this extension, the bELSM at the lowest order could reproduce the empirical spectra of the scalar mesons and nucleons, as well as the NM properties at saturation density.
Explicitly, we found that without introducing the 12-quark terms as in the previous work, the reasonable behavior of nuclear matter at high densities can be well reproduced. The couplings related to $\rho$ and $\omega$ mesons can be separated due to the new flavor structures, allowing the $E_{\rm sym}$ plateau to appear at the expected density range, around $2n_{0}$, instead of below $n_{0}$ as in the previous work. This is consistent with Ref. Li et al. (2022), which highlights the importance of such $E_{\rm sym}$ behavior to meet the constraints of GW170817 and the neutron skin thickness of ${}^{208}\rm Pb$ with a unified model. Besides the phenomena at densities, the pion decay constant is also consistent with recent lattice QCD results Aoki et al. (2020), with $f_{\pi}\approx 134~{}\rm MeV$ in el-g30eg, compared to $f_{\pi}\sim(60-70)~{}\rm MeV$ in the previous work.
With this improved lowest-order bELSM and the corresponding parameter space, a further study on the EOS over a wider density region was performed. Several parameter sets were found to meet the NM properties at $n_{0}$. The differences lie in the magnitude of $g_{aNN}$ and the four-vector meson coupling, $\tilde{g}_{3}$. It was found that small $g_{aNN}$ and $\tilde{g}_{3}$ lead to more compact NSs, consistent with the constraints of MSP J0740+6620 and GW170817. The influences of the EOS behaviors on NS M-R relations were also discussed: a soft $E_{\rm sym}$ at intermediate densities, with a stiff behavior at subsaturation densities, is necessary for the compactness of NSs to meet the GW170817 constraints; a stiff SV behavior at high densities is necessary for the $M_{\rm max}$ of NSs to meet the MSP J0740+6620 constraints. Besides the well-produced NS structures from el-g30eg and FSU-$\delta 6.7$, the SV behavior of el-g30eg is more favored, as its SV behavior can approach the conformal limit at high densities, expected for the fact that QCD is asymptoticly free at high energies.
The percentages of 2-quark and 4-quark constituents making up the scalar mesons were also calculated. The $\sigma$ meson was found to be composed of 2-quark and 4-quark constituents in similar proportions, while the $a_{0}$ meson is mainly composed of 4-quark constituents in the current optimal parameter spaces. This is consistent with previous analyses in vacuum Fariborz et al. (2008). The variation of the constituents due to different magnitudes of $g_{aNN}$ and four-vector meson couplings were also provided for future studies on scalar mesons and the macroscopic properties of compact objects in the universe. However, the quark constituents make the couplings related to $a_{0}$, both types: $g_{a_{0}NN}$ and $C_{\sigma^{i}\delta^{2}}$, different from the results of FSU-$\delta 6.7$ and Zabari-19, which are used for the symmetry energy at high densities. This again signifies the importance of considering the chiral symmetry pattern of QCD at low energies in phenomenological analyses beyond the saturation density.
In summary, the lowest-lying isoscalar and isovector scalar and vector mesons are all included within a unified model, bELSM, which also respects the chiral symmetry pattern of QCD at low energies. The NM properties and NS structures are systematically studied. Our recent works Guo et al. (2024a); Ma and Ma (2024); Guo et al. (2024b); Zhang et al. (2024a, b) focus on the relationships between the QCD symmetry pattern and compact objects in the universe. In the near future, improved AI analysis on a general Walecka-type model with the inclusion of all types of lowest-lying mesons and baryons will be available, which will help further investigate bELSM at densities due to its numerical complexity. Additionally, studies based on bELSM with the inclusion of strange freedoms are underway to address the properties of compact objects with hyperons. Meanwhile, the Hartree-Fock approach with $\chi\rm EFT$ will also be initially completed to study the EOS of NM, including quantum effects, which are ignored in the current RMF approximation.
Acknowledgements.
The work of Y. L. M. is supported in part by the National Science Foundation of China (NSFC) under Grant No. 12347103, the National Key R&D Program of China under Grant No. 2021YFC2202900 and Gusu Talent Innovation Program under Grant No. ZXL2024363.
Appendix A The lowest order bELSM
In this appendix, we provide details of the bELSM at the lowest order before the RMF approximation. In the model, the scalar mesons $S$ and pseudoscalar mesons $P$ are combined into a compact form $\Phi=S+iP$:
$\Phi=S+iP=\left(\begin{array}[]{ccc}\frac{\left(\sigma_{N}+a_{0}^{0}\right)+i\left(\eta_{N}+\pi^{0}\right)}{\sqrt{2}}&a_{0}^{+}+i\pi^{+}&K_{S}^{+}+iK^{+}\\ a_{0}^{-}+i\pi^{-}&\frac{\left(\sigma_{N}-a_{0}^{0}\right)+i\left(\eta_{N}-\pi^{0}\right)}{\sqrt{2}}&K_{S}^{0}+iK^{0}\\ K_{S}^{-}+iK^{-}&\bar{K}_{S}^{0}+i\bar{K}^{0}&\sigma_{S}+i\eta_{S}\end{array}\right)\ .$ (11)
The representations of the 2-quark and 4-quark (pseudo-)scalar configurations are similar, but with ”prime” and ”hat-prime” indices, as defined in the main text. For the vector and axial-vector mesons, they are expressed as:
$(R,L)^{\mu}=V^{\mu}\pm A^{\mu}=\frac{1}{\sqrt{2}}\left(\begin{array}[]{ccc}\frac{\omega_{N}^{\mu}+\rho^{\mu 0}}{\sqrt{2}}\pm\frac{f_{1N}^{\mu}+a_{1}^{\mu 0}}{\sqrt{2}}&\rho^{\mu+}\pm a_{1}^{\mu+}&K^{*\mu+}\pm K_{1}^{\mu+}\\ \rho^{\mu-}\pm a_{1}^{\mu-}&\frac{\omega_{N}^{\mu}-\rho^{\mu 0}}{\sqrt{2}}\pm\frac{f_{1N}^{\mu}-a_{1}^{\mu 0}}{}&K^{*\mu 0}\pm K_{1}^{\mu 0}\\ K^{*\mu-}\pm K_{1}^{\mu-}&\bar{K}^{*\mu 0}\pm\bar{K}_{1}^{\mu 0}&\omega_{S}^{\mu}\pm f_{1S}^{\mu}\end{array}\right)\ ,$ (12)
where $L(R)^{\mu}$ stands for the left(right)-handed vector mesons, and $V(A)^{\mu}$ denotes the vector(axial-vector) mesons. The baryon octet has expression:
$B=\left(\begin{array}[]{ccc}\frac{\Lambda}{\sqrt{6}}+\frac{\Sigma^{0}}{\sqrt{2}}&\Sigma^{+}&p\\ \Sigma^{-}&\frac{\Lambda}{\sqrt{6}}-\frac{\Sigma^{0}}{\sqrt{2}}&n\\ \Xi^{-}&\Xi^{0}&-\frac{2\Lambda}{\sqrt{6}}\end{array}\right)\ .$ (13)
The chiral symmetry transformations for (pseudo-)scalar configurations are given by Fariborz et al. (2005, 2008, 2009):
$\Phi^{\prime}\rightarrow g_{\mathrm{L}}\Phi^{\prime}g_{\mathrm{R}}^{\dagger},\quad\hat{\Phi}^{\prime}\rightarrow g_{\mathrm{L}}\hat{\Phi}^{\prime}g_{\mathrm{R}}^{\dagger}\ ,$ (14)
where $g_{\mathrm{L},\mathrm{R}}\in\mathrm{SU}(3)_{\mathrm{L},\mathrm{R}}$. They remain unchanged for $\rm U(1)_{V}$ transformation. But for $\rm U(1)_{A}$ transformation, they transform as:
$\Phi^{\prime}\rightarrow e^{2i\nu}\Phi^{\prime},\quad\hat{\Phi}^{\prime}\rightarrow e^{-4i\nu}\hat{\Phi}^{\prime}\ ,$ (15)
with $\nu$ being the chiral angle, and $q\bar{q}$ and $(qq)(\bar{q}\bar{q})$ configurations are assumed for 2-quark and 4-quark components, respectively. For the vector and axial-vector mesons, their chiral transformations are Parganlija et al. (2013):
$L_{\mu}\rightarrow g_{\mathrm{L}}L_{\mu}g_{\mathrm{L}}^{\dagger},\quad R_{\mu}\rightarrow g_{\mathrm{R}}R_{\mu}g_{\mathrm{R}}^{\dagger}\ ,$ (16)
with $\bar{q}_{\mathrm{L(R)}}\gamma_{\mu}q_{\mathrm{L(R)}}$ configuration assumed. They are both invariant for both $\rm U(1)_{V}$ and $\rm U(1)_{A}$ transformations.
In this work, the diquark approximation for baryons is adopted Olbrich (2015); Olbrich et al. (2016). The baryons are denoted as $N_{\mathrm{R}}^{(\mathrm{RR})}$, $N_{\mathrm{L}}^{(\mathrm{LL})}$, $N_{\mathrm{L}}^{(\mathrm{RR})}$, and $N_{\mathrm{R}}^{(\mathrm{LL})}$, where the superscripts represent the quark configuration and the subscripts represent the diquark configuration. The chiral transformations for baryons are:
$N_{\mathrm{R}}^{(\mathrm{RR})}$$\rightarrow$ $g_{\mathrm{R}}N_{\mathrm{R}}^{(\mathrm{RR})}g_{\mathrm{R}}^{\dagger}\ ,\quad N_{\mathrm{L}}^{(\mathrm{RR})}\rightarrow g_{\mathrm{L}}N_{\mathrm{L}}^{(\mathrm{RR})}g_{\mathrm{R}}^{\dagger}\ ,$
$N_{\mathrm{R}}^{(\mathrm{LL})}$$\rightarrow$ $g_{\mathrm{R}}N_{\mathrm{R}}^{(\mathrm{LL})}g_{\mathrm{L}}^{\dagger}\ ,\quad N_{\mathrm{L}}^{(\mathrm{LL})}\rightarrow g_{\mathrm{L}}N_{\mathrm{L}}^{(\mathrm{LL})}g_{\mathrm{L}}^{\dagger}\ .$ (17)
And the $\rm U(1)_{A}$ transformations are:
$N_{\mathrm{R}}^{(\mathrm{RR})}$$\rightarrow$ $e^{-3iv}N_{\mathrm{R}}^{(\mathrm{RR})}\ ,\quad N_{\mathrm{L}}^{(\mathrm{RR})}\rightarrow e^{-iv}N_{\mathrm{L}}^{(\mathrm{RR})}\ ,$
$N_{\mathrm{R}}^{(\mathrm{LL})}$$\rightarrow$ $e^{iv}N_{\mathrm{R}}^{(\mathrm{LL})}\ ,\quad N_{\mathrm{L}}^{(\mathrm{LL})}\rightarrow e^{3iv}N_{\mathrm{L}}^{(\mathrm{LL})}\ ,$ (18)
with invariant under $\rm U(1)_{V}$ transformation. The physical baryons are defined as:
$N_{\mathrm{R},\mathrm{~{}L}}^{(\mathrm{RR})}=\frac{1}{\sqrt{2}}\frac{1\pm\gamma_{5}}{2}B\ ,\quad N_{\mathrm{R},\mathrm{~{}L}}^{(\mathrm{LL})}=-\frac{1}{\sqrt{2}}\frac{1\pm\gamma_{5}}{2}B\ ,$ (19)
with excited state $N^{*}$ being neglected since it is irrelevant to the present work.
Before constructing the full Lagrangian, a series of power-counting rules are needed to organize the effective operators:
  • The operators are limited to dimension-4, as higher-dimensional operators are suppressed by the cutoff scale.
  • The quark number of an operator is limited to 8 Fariborz et al. (2005, 2008, 2009), as operators with more quarks are suppressed in the effective description of quark lines.
  • The number of flavor space traces is limited to 1, since more trace terms are suppression by $N_{c}$ Parganlija et al. (2010).
  • The explicit symmetry breaking caused by quark mass is treated as a perturbation and is ignored in the current work.
Consequently, the full Lagrangian at the lowest order is
$\mathcal{L}_{\rm LO}=\mathcal{L}_{\rm M}+\mathcal{L}_{\rm V}+\mathcal{L}_{\rm B}.$ (20)
The (pseudo-)scalar meson part is Fariborz et al. (2008):
$\mathcal{L}_{\rm M}$ $=$ $\frac{1}{2}\operatorname{Tr}\left(\partial_{\mu}\Phi^{\prime}\partial^{\mu}\Phi^{\prime\dagger}\right)+\frac{1}{2}\operatorname{Tr}\left(\partial_{\mu}\hat{\Phi}^{\prime}\partial^{\mu}\hat{\Phi}^{\prime\dagger}\right)+c_{2}\operatorname{Tr}\left(\Phi^{\prime}\Phi^{\prime\dagger}\right)-c_{4}\operatorname{Tr}\left(\Phi^{\prime}\Phi^{\prime\dagger}\Phi^{\prime}\Phi^{\prime\dagger}\right)-d_{2}\operatorname{Tr}\left(\hat{\Phi}^{\prime}\hat{\Phi}^{\prime\dagger}\right)$ (21)
${}-e_{3}\left(\epsilon_{abc}\epsilon^{def}\Phi_{d}^{\prime a}\Phi_{e}^{\prime b}\hat{\Phi}_{f}^{\prime c}+{\rm h.c.}\right)-c_{3}\left[\gamma_{1}\ln\left(\frac{\operatorname{det}\Phi^{\prime}}{\operatorname{det}{\Phi}^{\prime\dagger}}\right)+\left(1-\gamma_{1}\right)\ln\left(\frac{\operatorname{Tr}\left(\Phi^{\prime}\hat{\Phi}^{\prime\dagger}\right)}{\operatorname{Tr}\left(\hat{\Phi}^{\prime}\Phi^{\prime\dagger}\right)}\right)\right]^{2}\ ,$
where the $c_{3}$ term accounts for $\rm U(1)_{A}$ anomaly at low energies. The (axial-)vector meson part is written as:
$\mathcal{L}_{\rm V}$ $=$ ${}-\frac{1}{8}{\rm Tr}\left({L}_{\mu\nu}{L}^{\mu\nu}+{R}_{\mu\nu}{R}^{\mu\nu}\right)$ (22)
${}+g_{1}\left[{\rm Tr}\left(\partial_{\nu}{R}_{\mu}{R}^{\mu}{R}^{\nu}\right)+{\rm Tr}\left(\partial_{\nu}{L}_{\mu}{L}^{\mu}{L}^{\nu}\right)\right]+g_{2}\left[{\rm Tr}\left(\partial_{\nu}{R}_{\mu}{R}^{\nu}{R}^{\mu}\right)+{\rm Tr}\left(\partial_{\nu}{L}_{\mu}{L}^{\nu}{L}^{\mu}\right)\right]$
${}+g_{3}\left[{\rm Tr}\left({R}_{\mu}{R}^{\mu}{R}_{\nu}{R}^{\nu}\right)+{\rm Tr}\left({L}_{\mu}{L}^{\mu}{L}_{\nu}{L}^{\nu}\right)\right]+g_{4}\left[{\rm Tr}\left({R}_{\mu}{R}^{\nu}{R}_{\mu}{R}^{\nu}\right)+{\rm Tr}\left({L}_{\mu}{L}^{\nu}{L}_{\mu}{L}^{\nu}\right)\right]$
${}+h_{1}\left[{\rm Tr}\left(\Phi^{\prime\dagger}\Phi^{\prime}R_{\mu}R^{\mu}\right)+{\rm Tr}\left(\Phi^{\prime}\Phi^{\prime\dagger}L_{\mu}L^{\mu}\right)\right]+h_{2}\left[{\rm Tr}\left(L^{\mu}\Phi^{\prime}R_{\mu}\Phi^{\prime\dagger}\right)\right]+h_{3}\left[{\rm Tr}\left(\Phi^{\prime\dagger}\partial_{\mu}\Phi^{\prime}R^{\mu}\right)+{\rm Tr}\left(\Phi^{\prime}\partial_{\mu}\Phi^{\prime\dagger}L^{\mu}\right)\right]$
${}+\frac{a_{1}}{2}\epsilon_{abc}\epsilon^{def}\left[\left(R_{\mu}\right)_{ad}\left(R_{\nu}\right)_{be}\left(\Phi^{\prime\dagger}\Phi\right)_{cf}+\left(L_{\mu}\right)_{ad}\left(L_{\nu}\right)_{be}\left(\Phi\Phi^{\prime\dagger}\right)_{cf}\right]$
${}+\frac{a_{2}}{2}\epsilon_{abc}\epsilon^{def}\left[\left(R_{\mu}\right)_{ad}\left(R_{\nu}\right)_{be}\left(R^{\mu}R^{\nu}\right)_{cf}+\left(L_{\mu}\right)_{ad}\left(L_{\nu}\right)_{be}\left(L^{\mu}L^{\nu}\right)_{cf}\right]$
${}+\frac{a_{3}}{2}\epsilon_{abc}\epsilon^{def}\left[\left(R_{\mu}\right)_{ad}\left(R^{\mu}\right)_{be}\left(R^{\nu}R_{\nu}\right)_{cf}+\left(L_{\mu}\right)_{ad}\left(L^{\mu}\right)_{be}\left(L^{\nu}L_{\nu}\right)_{cf}\right]$
${}+\frac{a_{4}}{2}\epsilon_{abc}\epsilon^{def}\left[\left(R_{\mu}\right)_{ad}\left(R^{\nu}\right)_{be}\left(\partial^{\mu}R_{\nu}\right)_{cf}+\left(L_{\mu}\right)_{ad}\left(L^{\nu}\right)_{be}\left(\partial^{\mu}L_{\nu}\right)_{cf}\right]\ .$
The baryon part has expression
$\mathcal{L}_{\rm B}$ $=$ ${\rm Tr}\left(\bar{N}_{\rm R}^{\left(\rm RR\right)}i\not{\partial}N_{\rm R}^{\left(\rm RR\right)}+\bar{N}_{\rm L}^{\left(\rm RR\right)}i\not{\partial}N_{\rm L}^{\left(\rm RR\right)}+\bar{N}_{\rm R}^{\left(\rm LL\right)}i\not{\partial}N_{\rm R}^{\left(\rm LL\right)}+\bar{N}_{\rm L}^{\left(\rm LL\right)}i\not{\partial}N_{\rm L}^{\left(\rm LL\right)}\right)$ (23)
[…]
${}+n_{3}{\rm Tr}\left(\bar{N}_{\rm R}^{\left(\rm RR\right)}\gamma_{\mu}N_{\rm R}^{\left(\rm RR\right)}R^{\mu}+\bar{N}_{\rm L}^{\left(\rm LL\right)}\gamma_{\mu}N_{\rm L}^{\left(\rm RR\right)}L^{\mu}\right)+n_{4}\left(\bar{N}_{\rm R}^{\left(\rm LL\right)}\gamma_{\mu}N_{\rm R}^{\left(\rm LL\right)}L^{\mu}+\bar{N}_{\rm L}^{\left(\rm RR\right)}\gamma_{\mu}N_{\rm L}^{\left(\rm RR\right)}R^{\mu}\right)$
${}+n_{a}\epsilon_{abc}\epsilon^{def}\left[\left(\bar{N}_{\rm R}^{\left(\rm RR\right)}\right)_{ad}\gamma_{\mu}\left(N_{\rm R}^{\left(\rm RR\right)}\right)_{be}\left(R^{\mu}\right)_{cf}+\left(\bar{N}_{\rm L}^{\left(\rm LL\right)}\right)_{ad}\gamma_{\mu}\left(N_{\rm L}^{\left(\rm LL\right)}\right)_{be}\left(L^{\mu}\right)_{cf}\right]$
${}+n_{g}{\rm Tr}\left[\bar{N}_{\rm R}^{\left(\rm RR\right)}\Phi^{\prime\dagger}N_{\rm L}^{\left(\rm RR\right)}+\bar{N}_{\rm L}^{\left(\rm LL\right)}\Phi^{\prime}N_{\rm R}^{\left(\rm LL\right)}+{\rm h.c.}\right]$
${}+n_{e}\epsilon_{abc}\epsilon^{def}\left[\left(\bar{N}_{\rm R}^{\left(\rm RR\right)}\right)_{ad}\left(\Phi^{\prime\dagger}\right)_{be}\left(N_{\rm L}^{\left(\rm RR\right)}\right)_{cf}+\left(\bar{N}_{\rm L}^{\left(\rm LL\right)}\right)_{ad}\left(\Phi^{\prime}\right)_{be}\left(N_{\rm R}^{\left(\rm LL\right)}\right)_{cf}+{\rm h.c.}\right]\ .$
And after substituting Eq. (19) into the Lagrangian, the baryon part can be rewritten as
$\mathcal{L}_{\rm B}$ $=$ ${\rm Tr}\left(\bar{B}i\not{\partial}B\right)+c{\rm Tr}\left(\bar{B}\gamma_{\mu}V^{\mu}B+\bar{B}\gamma_{\mu}\gamma_{5}A^{\mu}B\right)+c^{\prime}{\rm Tr}\left(\bar{B}\gamma_{\mu}BV^{\mu}+\bar{B}\gamma_{\mu}\gamma_{5}BA^{\mu}\right)$ (24)
${}+h\epsilon_{abc}\epsilon^{def}\left[\left(\bar{B}\right)_{ad}\gamma_{\mu}\left(B\right)_{be}\left(V^{\mu}\right)_{cf}+\left(\bar{B}\right)_{ad}\gamma_{\mu}\gamma_{5}\left(B\right)_{be}\left(A^{\mu}\right)_{cf}\right]$
${}-\frac{g}{2}{\rm Tr}\left[\bar{B}\left(\Phi^{\prime}+\Phi^{\prime\dagger}\right)B+\bar{B}\gamma_{5}\left(\Phi^{\prime}-\Phi^{\prime\dagger}\right)B\right]$
${}-\frac{e}{2}\epsilon_{abc}\epsilon^{def}\left[\left(\bar{B}\right)_{ad}\left(\Phi^{\prime}+\Phi^{\prime\dagger}\right)_{be}\left(B\right)_{cf}+\left(\bar{B}\right)_{ad}\gamma_{5}\left(\Phi^{\prime}-\Phi^{\prime\dagger}\right)_{be}\left(B\right)_{cf}\right]\ ,$
where $c=(c_{1}+c_{2})/2$, $c^{\prime}=(c_{3}+c_{4})/2$, $h=n_{a}/2$, $g=-n_{g}$ and $e=-n_{e}$.
Appendix B Approximations in the current work
In the present study of NM properties, we take RMF, which is a classical approximation, where meson fields are treated as potentials induced by matter fields, such as nucleons in the current work. Additionally, the system composed of these baryons is considered to be infinite, homogeneous, and static, so that any quantum numbers and currents are averaged to zero. This results in the following approximations
$\langle\omega^{\mu}\rangle$ $=$ $\langle\omega^{0}\rangle\delta^{\mu,0}=\omega\delta^{\mu,0}\ ,$
$\langle\rho_{i}^{\mu}\rangle$ $=$ $\langle\rho_{0}^{0}\rangle\delta_{i,0}\delta^{\mu,0}=\rho\delta_{i,0}\delta^{\mu,0}\ ,$
$\langle\partial_{\mu}\sigma\rangle$ $=$ $\partial_{\mu}\langle\sigma\rangle=0\ ,$
$\langle\pi^{i}\rangle$ $=$ $0\ ,$ (25)
and other related meson fields are also treated similarly. For instance, $K$ mesons are excluded due to their non-zero strange quantum number.
Mesons heavier than $1\rm GeV$, such as $a^{\prime}_{0}$ and $\omega_{S}$, are ignored since their contributions are suppressed at the current working scale. This leads to the following parameter choices in the current work: $\tilde{g}_{3}=2(g_{3}+g_{4})$, $\tilde{h}_{2}=2h_{1}+h_{2}$, while the other parameters remain the same as the original ones after the approximations. Details on how to solve the RMF equations of motion (EOMs) to obtain the equation of state (EOS) after these approximations are provided in our other work Guo et al. (2024b).

### v3

@@
 Abstract
-It has been argued that the iso-scalar and iso-vector mesons play significant roles in nuclear matter and neutron star structures.
+It has been argued that the isoscalar and isovector mesons play significant roles in nuclear matter and neutron star structures.
 We improve the extended linear sigma model with baryons, proposed in our previous work, by introducing the flavor structures constructed from antisymmetric tensors of chiral representations to study these physics.
-The parameter space of this model is refined with well-reproduced nuclear matter properties at saturation density by the lowest order Lagrangian, ensuring consistency with vacuum results, such as $f_{\pi}\approx 134\,\text{MeV}$.
-The anticipated plateau-like behaviors of the symmetry energy are predicted at intermediate densities, which is crucial for the consistency of GW170817 and the neutron skin thickness of $\text{Pb}^{208}$.
+The parameter space of this model is refined with well-reproduced nuclear matter properties at saturation density by the lowest-order Lagrangian, ensuring consistency with vacuum results, such as $f_{\pi}\approx 134\,\text{MeV}$.
+The anticipated plateaulike behaviors of the symmetry energy are predicted at intermediate densities, which is crucial for the consistency of GW170817 and the neutron skin thickness of $\text{Pb}^{208}$.
 Subsequently, neutron star structures are calculated using several parameter sets, and the results for the nuclear matter properties at saturation density align with empirical values.
-It is found that the neutron star structures are sensitive to the couplings between the iso-vector $a_{0}$ meson and nucleons and the four-vector meson couplings: small values of both are favorable.
+It is found that the neutron star structures are sensitive to the couplings between the isovector $a_{0}$ meson and nucleons and the four-vector meson couplings: small values of both are favorable.
 Meanwhile, nuclear matter properties at saturation density favor larger values of the latter and are not sensitive to the former.
@@
 I Introduction
-Nuclear matter (NM), a system dominated by the nonperturbative quantum chromodynamics (QCD), has been studied for several decades.
+Nuclear matter (NM), a system dominated by the nonperturbative QCD, has been studied for several decades.
 But there are still great uncertainties about its properties, especially at compact star densities, due to our little knowledge on the phase structure of strong interaction matter Fukushima and Hatsuda (2011); Lattimer and Prakash (2016); Chen (2017); Baym et al.
@@
 In modeling NM, based on the phenomenological one-boson exchange for nucleon interactions Holinde and Machleidt (1975); Erkelenz (1974); Nagels et al.
-(1978); Machleidt (1989), it has long been recognized that the scalar meson resonances, such as $\sigma,f_{0}$ and $a_{0}(\delta)$, play indispensable roles by using the mean field approach (MFA) Walecka (1974); Serot and Walecka (1986, 1997) [^1] Hereafter, following the Particle Data Group, we denote the iso-vector scalar mesons as $a_{0}$.
-Without the iso-scalar scalar mesons $\sigma/f_{0}$, which provide the attractive force between nucleons, it’s hard to obtain the empirical values of the saturation density $n_{0}\simeq 0.16$ fm-3 and binding energy $E_{0}\simeq 16.0$ MeV Holinde and Machleidt (1975); Erkelenz (1974); Nagels et al.
+(1978); Machleidt (1989), it has long been recognized that the scalar meson resonances, such as $\sigma,f_{0}$ and $a_{0}$, play indispensable roles by using the mean field approach (MFA) Walecka (1974); Serot and Walecka (1986, 1997) [^1] Hereafter, following the Particle Data Group, we denote the isovector scalar mesons as $a_{0}$.
+Without the isoscalar scalar mesons $\sigma/f_{0}$, which provide the attractive force between nucleons, it’s hard to obtain the empirical values of the saturation density $n_{0}\simeq 0.16$ fm-3 and binding energy $E_{0}\simeq 16.0$ MeV Holinde and Machleidt (1975); Erkelenz (1974); Nagels et al.
 (1978); Machleidt (1989).
-And the self-interactions of $\sigma$ meson is necessary to reduce the nuclear incompressibility coefficient of symmetric nuclear matter Boguta and Bodmer (1977) and other empirical values Motohiro et al.
+And the self-interactions of the $\sigma$ meson are necessary to reduce the nuclear incompressibility coefficient of symmetric nuclear matter Boguta and Bodmer (1977) and other empirical values Motohiro et al.
 (2015).
-In addition to the iso-scalar scalar mesons, the iso-vector scalar mesons are important for describing the asymmetric nuclear matter and hence the properties of neutron stars Kubis and Kutschera (1997); Hofmann et al.
+In addition to the isoscalar scalar mesons, the isovector scalar mesons are important for describing the asymmetric nuclear matter and hence the properties of neutron stars Kubis and Kutschera (1997); Hofmann et al.
 (2001); Liu et al.
@@
 (2025).
-Furthermore, it is found that the scalar-iso-scalar scalar meson couplings, e.g., $\sigma a_{0}^{2}$ and $\sigma^{2}a_{0}^{2}$ have large effects on both the symmetry energy $E_{\rm sym}$ and its slope $L$ Zabari et al.
+Furthermore, it is found that the scalar-isoscalar scalar meson couplings, e.g., $\sigma a_{0}^{2}$ and $\sigma^{2}a_{0}^{2}$ have large effects on both the symmetry energy $E_{\rm sym}$ and its slope $L$ Zabari et al.
 (2019); Kubis et al.
@@
 Normally, in the models of NM, the scalar mesons are included with respect to that quantum numbers, such as parity, spin and isospin, less attention has been paid to other properties related to QCD.
-From QCD point of view, the structures of scalar mesons are still under debate, especially the light ones (see, e.g., Close and Tornqvist (2002) for a review).
+From a QCD point of view, the structures of scalar mesons are still under debate, especially the light ones (see, e.g., Ref.
+Close and Tornqvist (2002) for a review).
 From the mass ordering of the scalar mesons below $1$ GeV, it is difficult to regard them as the pure two-quark states but an admixture of the two- and four-quark configurations.
-Considering that the two quark and four quark components of scalar mesons have different axial transformation properties, we recently proposed an extended linear sigma model (ELSM) with baryons, taken as diquark-quark configurations, (bELSM) to study the properties of NM Ma and Ma (2024).
+Considering that the two-quark and four-quark components of scalar mesons have different axial transformation properties, we recently proposed an extended linear sigma model (ELSM) with baryons, where baryons are taken as diquark-quark configurations, to study the properties of NM, and the model is noted as baryonic extended linear sigma model (bELSM).
+Ma and Ma (2024).
 In bELSM, the masses of the mesons and baryons are generated from the spontaneous chiral symmetry breaking.
-Due to the conservation of axial transformation, only the 2-quark component of the scalar mesons couples to the baryons at the lowest order, which may provide deeper insights into the constituents of the scalar mesons from the NM and NS properties.
+Because of the conservation of axial transformation, only the two-quark component of the scalar mesons couples to the baryons at the lowest-order, which may provide deeper insights into the constituents of the scalar mesons from the NM and NS properties.
 The $\rm U(1)_{A}$ anomaly is involved in the ’t Hooft determinant Rosenzweig et al.
@@
 (2016).
-After pinning down the NM properties at saturation density $n_{0}$, a plateau-like structure of the symmetry energy, $E_{\rm sym}$, at intermediate densities, $\sim 2n_{0}$, was found in the bELSM.
+After pinning down the NM properties at saturation density $n_{0}$, a plateaulike structure of the symmetry energy, $E_{\rm sym}$, at intermediate densities, $\sim 2n_{0}$, was found in the bELSM.
 This is consistent with the analysis based on various of approaches in Refs.
@@
 (2016).
-We also investigated the effects of parameter variations on different coefficients of Taylor expansions of energy per-nucleon, $E=\mathcal{E}/n$, by density, $n=n_{n}+n_{p}$, and asymmetry, $\alpha=(n_{n}-n_{p})/n$, to provide further guidance on the interaction parameterization of NM:
-$E(n,\alpha)$ $\simeq$ $E(n_{0},0)+E_{\rm sym}(n_{0})\alpha^{2}+L(n_{0})\alpha^{2}\chi+\frac{K(n_{0})}{2!}\chi^{2}+\frac{J(n_{0})}{3!}\chi^{3}+\mathcal{O}\left(\alpha^{4}\right)+\mathcal{O}\left(\chi^{4}\right)+\mathcal{O}\left(\alpha^{2}\chi^{2}\right)$ (1)
-$\simeq$ $E_{0}+\left.\frac{\alpha^{2}}{2}\frac{\partial^{2}E(n_{0},\alpha)}{\partial\alpha^{2}}\right|_{\alpha=0}+\alpha^{2}\chi\left.\left(3n\frac{\partial E_{\mathrm{sym}}(n)}{\partial n}\right)\right|_{n=n_{0}}+\frac{\chi^{2}}{2!}\left.\left(9n^{2}\frac{\partial^{2}E_{0}(n)}{\partial n^{2}}\right)\right|_{n=n_{0}}$
-${}+\frac{\chi^{3}}{3!}\left.\left(27n^{3}\frac{\partial^{3}E_{0}(n)}{\partial n^{3}}\right)\right|_{n=n_{0}}+\mathcal{O}\left(\alpha^{4}\right)+\mathcal{O}\left(\chi^{4}\right)+\mathcal{O}\left(\alpha^{2}\chi^{2}\right)\ ,$
+We also investigated the effects of parameter variations on different coefficients of Taylor expansions of energy per nucleon, $E=\mathcal{E}/n$, by density, $n=n_{n}+n_{p}$, and asymmetry, ${\delta}=(n_{n}-n_{p})/n$, to provide further guidance on the interaction parameterization of NM,
+$E(n,{\delta})$ $\simeq$ $E(n_{0},0)+E_{\rm sym}(n_{0}){\delta}^{2}+L(n_{0}){\delta}^{2}\chi+\frac{K(n_{0})}{2!}\chi^{2}+\frac{J(n_{0})}{3!}\chi^{3}+\mathcal{O}\left({\delta}^{4}\right)+\mathcal{O}\left(\chi^{4}\right)+\mathcal{O}\left({\delta}^{2}\chi^{2}\right)$ (1)
+$\simeq$ $E_{0}+\left.\frac{{\delta}^{2}}{2}\frac{\partial^{2}E(n_{0},{\delta})}{\partial{\delta}^{2}}\right|_{{\delta}=0}+{\delta}^{2}\chi\left.\left(3n\frac{\partial E_{\mathrm{sym}}(n)}{\partial n}\right)\right|_{n=n_{0}}+\frac{\chi^{2}}{2!}\left.\left(9n^{2}\frac{\partial^{2}E_{0}(n)}{\partial n^{2}}\right)\right|_{n=n_{0}}$
+${}+\frac{\chi^{3}}{3!}\left.\left(27n^{3}\frac{\partial^{3}E_{0}(n)}{\partial n^{3}}\right)\right|_{n=n_{0}}+\mathcal{O}\left({\delta}^{4}\right)+\mathcal{O}\left(\chi^{4}\right)+\mathcal{O}\left({\delta}^{2}\chi^{2}\right)\ ,$
 where $\chi\equiv\left(n-n_{0}\right)/3n_{0}$. $L(n_{0})$, $K(n_{0})$, and $J(n_{0})$ refer to the symmetry energy slope, incompressibility, and skewness of NM at $n_{0}$, respectively.
-However, when calculating the $E_{\rm sym}(n)$ using bELSM, the plateau-like structure is found to be lower than the results in Ref.
+However, when calculating the $E_{\rm sym}(n)$ using bELSM, the plateaulike structure is found to be lower than the results in Refs.
 Miyatsu et al.
@@
 (2022), leading to a loose structure of NS with tidal deformation $\Lambda_{1.4}>1000$.
-This is due to the couplings related to $\rho$ and $\omega$ mesons being constrained by the $\rm U(3)_{V}$ symmetry, which is not entirely realistic, especially in nuclear medium Paeng et al.
+This is due to the couplings related to $\rho$ and $\omega$ mesons being constrained by the $\rm U(3)_{V}$ symmetry, which is not entirely realistic, especially in the nuclear medium Paeng et al.
 (2012).
@@
 (1998), which can split the couplings of the singlet and octet vector mesons to baryons.
-In this work, we extend the bELSM by including this kind of operators in both the mesonic and baryonic sectors at the lowest order and improve the parameter space of the original bELSM.
-The 12-quark terms, which were necessary to maintain reasonable behavior at high densities in the work of Ref.
-Ma and Ma (2024), can be removed, enhancing the rationality of the power-counting in the bELSM.
-The vacuum expectation values (VEVs) of the scalar meson fields are close to the results of Ref.
+In this work, we extend the bELSM by including this kind of operators in both the mesonic and baryonic sectors at the lowest-order and improve the parameter space of the original bELSM.
+The 1two-quark terms, which were necessary to maintain reasonable behavior at high densities in the work of Ref.
+Ma and Ma (2024), can be removed, enhancing the rationality of the power counting in the bELSM.
+The vacuum expectation values of the scalar meson fields are close to the results of Ref.
 Fariborz et al.
 (2008), leading to a reasonable $f_{\pi}$ (quark condensate) and low-energy pion scattering results.
-With the improved bELSM at the lowest order, the NM properties at $n_{0}$ and the bare masses of the mesons and baryons can be well reproduced.
+With the improved bELSM at the lowest-order, the NM properties at $n_{0}$ and the bare masses of the mesons and baryons can be well reproduced.
 The bELSM is then applied to calculating NS structures.
@@
 This $K(n_{0})$ also results in an interesting behavior of SV, approaching the conformal limit at high densities, $n\sim 5n_{0}$, whereas other SVs from smaller $K(n_{0})$ will be too soft.
-This highlights the statistical significance of NS observations when obtaining realistic chiral effective field theories ($\chi$EFTs) or models Guo et al.
+This highlights the statistical significance of NS observations when obtaining realistic chiral effective field theories or models Guo et al.
 (2024a).
-In addition, the influence of portions in the 2-quark and 4-quark constituents of the scalar mesons on NS structures and NM properties is also shown and their physical magnitudes are estimated.
+In addition, the influence of portions in the two-quark and four-quark constituents of the scalar mesons on NS structures and NM properties is also shown and their physical magnitudes are estimated.
 These discussions set up a relationship between microscopic symmetry and macroscopic phenomena.
 The rest of this article is organized as follows: In Sec.
-II, we first introduce the improved bELSM at the lowest order under the relativistic mean field (RMF) approximation.
+II, we first introduce the improved bELSM at the lowest-order under the relativistic mean field (RMF) approximation.
 Then we study the properties of the scalar mesons on the NM properties and NS structures in detail.
@@
 III, the discussion on the results of the current work and the outlook are provided.
-The details of the model construction and approximation are given in Appendix.
-II Extended linear sigma model at densities
+The details of the model construction and approximation are given in the Appendix.
+II Extended linear sigma model at finite densities
 II.1 Theoretical framework
 In this work, we consider the lowest-order bELSM given in Ref.
-Ma and Ma (2024) including terms up to 8-quark lines with a single trace and operators of dimension-4.
+Ma and Ma (2024) including terms up to eight quark lines with a single trace and operators of dimension 4.
 In addition, as mentioned in the Introduction, we consider the terms with three-rank Levi-Civita tensor in flavor space.
@@
 ${}+a_{1}\epsilon_{abc}\epsilon_{def}V_{ad}V_{be}\left(S^{\prime 2}\right)_{cf}\ .$
-The $h,\ e$ and $a_{1}$ terms are the terms which were not considered in Ref. Ma and Ma (2024). The details of the model construction are given in App. A.
+The h,eh,\ e and $a_{1}$ terms are the terms which were not considered in Ref. Ma and Ma (2024). The details of the model construction are given in Appendix A.
 In Lagrangian (2), $S^{\prime}$ is the two-quark configuration of scalar states which can be decomposed as
 $S^{\prime}$ $=$ $\frac{1}{\sqrt{2}}(\lambda_{8}f_{0}^{\prime}+\lambda_{3}a_{0}^{\prime})+\frac{1}{\sqrt{3}}I\sigma^{\prime},$ (4)
-with $\lambda_{3}$ and $\lambda_{8}$ being the 3rd and 8th components of Gell-Mann matrices, and $I$ as the identity matrix.
+with $\lambda_{3}$ and $\lambda_{8}$ being the third and eighth components of Gell-Mann matrices and $I$ being the identity matrix.
 The similar decomposition applies to the four-quark configuration $\hat{S}^{\prime}$.
@@
 $B\rightarrow\left(\begin{array}[]{ccc}0&0&p\\ 0&0&n\\ 0&0&0\end{array}\right)\ .$ (6)
-So, with $p$ and $n$, there are iso-scalar scalar configurations $\sigma^{\prime}/\hat{\sigma}^{\prime}$, $f_{0}^{\prime}/\hat{f}_{0}^{\prime}$ and iso-vector scalar configurations $a_{0}^{\prime}/\hat{a}_{0}^{\prime}$ in model (2).
+So, with $p$ and $n$, there are isoscalar scalar configurations $\sigma^{\prime}/\hat{\sigma}^{\prime}$, $f_{0}^{\prime}/\hat{f}_{0}^{\prime}$ and isovector scalar configurations $a_{0}^{\prime}/\hat{a}_{0}^{\prime}$ in model (2).
 After chiral symmetry breaking, the mixing of these scalar configurations yields the physical states of scalar mesons.
-The spontaneous chiral symmetry breaking is achieved by the vacuum expectation values of both 2-quark and 4-quark fields: $\langle\sigma^{\prime}\rangle=\sqrt{3}\alpha$ and $\langle\hat{\sigma}^{\prime}\rangle=\sqrt{3}\beta$, which are determined by the minimum of the potential $\mathcal{L}_{\rm M}$:
-$\left\langle\frac{\partial V_{\rm M}}{\partial\sigma^{\prime}}\right\rangle$ $=$ $2\alpha\left(-c_{2}+2c_{4}\alpha^{2}+4e_{3}\beta\right)=0\ ,$
-$\left\langle\frac{\partial V_{\rm M}}{\partial\hat{\sigma}^{\prime}}\right\rangle$ $=$ $2\left(d_{2}\beta+2e_{3}\alpha^{2}\right)=0\ ,$ (7)
-where the three-flavor symmetry is used. The physical states $\sigma$, $a_{0}$, and $f_{0}$ are obtained by the mixing of the 2-quark and 4-quark configurations, via
+The spontaneous chiral symmetry breaking is achieved by the vacuum expectation values of both two-quark and four-quark fields: $\langle\sigma^{\prime}\rangle=\sqrt{3}\alpha$ and $\langle\hat{\sigma}^{\prime}\rangle=\sqrt{3}\beta$, which are determined by the minimum of the potential $\mathcal{L}_{\rm M}$,
+$\left\langle\frac{\partial\mathcal{L}_{\rm M}}{\partial\sigma^{\prime}}\right\rangle$ $=$ $2\alpha\left(-c_{2}+2c_{4}\alpha^{2}+4e_{3}\beta\right)=0\ ,$
+$\left\langle\frac{\partial\mathcal{L}_{\rm M}}{\partial\hat{\sigma}^{\prime}}\right\rangle$ $=$ $2\left(d_{2}\beta+2e_{3}\alpha^{2}\right)=0\ ,$ (7)
+where the three-flavor symmetry is used. The physical states $\sigma$, $a_{0}$, and $f_{0}$ are obtained by the mixing of the two-quark and four-quark configurations, via
 $\sigma=\cos\theta_{0}\sigma^{\prime}+\sin\theta_{0}\hat{\sigma}^{\prime}\ ,$
@@
 The physical quantities that can be used to constrain these parameters are the bare masses of the scalar mesons and nucleons, as well as the NM properties around saturation density, $n_{0}$.
-The empirical constraints and corresponding fitted values are shown in Tabs.
-2 and 3, and the fitted parameters are listed in Tab.
-1.
+The empirical constraints and corresponding fitted values are shown in Tables 2 and 3, and the fitted parameters are listed in Table 1.
 Table 1: Values of the parameters used in this work.
@@
 The other parameters are dimensionless.
-The parameter sets are denoted as el-g30g, el-g30e, el-g30eg, el-g350eg, el-g3100eg, and el-g3150eg, respectively.
-’el’ refers to the bELSM; ’g3’ refers to the four-vector meson coupling in Eq.
-(3c), with the number following it denoting the corresponding magnitude; ’e’ or ’g’ indicates that the coupling between baryons and scalar mesons is mainly from ’e’ or ’g’ terms in Eq.
-(3a), and ’eg’ indicates that the contributions of the ’e’ and ’g’ terms are similar.
+The parameter sets are denoted as el-g30g, el-g30e, el-g30eg, el-g350eg, el-g3100eg, and el-g3150eg, respectively; ”el” refers to the bELSM; ”g3” refers to the four-vector meson coupling in Eq.
+(3c), with the number following it denoting the corresponding magnitude; ”e” or ”g” indicates that the coupling between baryons and scalar mesons is mainly from ”e” or ”g” terms in Eq.
+(3a), and ”eg” indicates that the contributions of the ”e” and ”g” terms are similar.
 $\alpha$ $\beta$ $e_{3}$ $c_{4}$ $\tilde{h}_{2}$ $\tilde{g}_{3}$ $c$ $g$ $a_{1}$ $h$ $e$
-el-g30g $61.1$ $24.7$ $-2.06$ $44.0$ $80.1$ $1.59$ $-0.792$ $15.4$ $4.25$ $11.4$ $-0.0270$
-el-g30e $61.1$ $24.4$ $-2.05$ $43.6$ $80.0$ $0.542$ $11.4$ $0.234$ $4.17$ $-0.790$ $15.1$
-el-g30eg $61.4$ $26.4$ $-2.10$ $45.6$ $79.3$ $0.397$ $9.51$ $6.54$ $4.10$ $-2.61$ $8.75$
-el-g350eg $61.2$ $25.6$ $-2.10$ $44.4$ $79.9$ $51.5$ $10.1$ $6.35$ $4.14$ $-2.65$ $9.00$
-el-g3100eg $60.8$ $24.0$ $-2.09$ $42.4$ $80.8$ $100$ $10.6$ $7.10$ $4.19$ $-2.88$ $8.34$
-el-g3150eg $60.7$ $24.3$ $-2.11$ $42.0$ $81.0$ $150$ $11.1$ $7.15$ $4.18$ $-3.05$ $8.30$
+el-g30g $61.1$ $24.7$ 2.06-2.06 $44.0$ $80.1$ $1.59$ 0.792-0.792 $15.4$ $4.25$ $11.4$ 0.0270-0.0270
+el-g30e $61.1$ $24.4$ 2.05-2.05 $43.6$ $80.0$ $0.542$ $11.4$ $0.234$ $4.17$ 0.790-0.790 $15.1$
+el-g30eg $61.4$ $26.4$ 2.10-2.10 $45.6$ $79.3$ $0.397$ $9.51$ $6.54$ $4.10$ 2.61-2.61 $8.75$
+el-g350eg $61.2$ $25.6$ 2.10-2.10 $44.4$ $79.9$ $51.5$ $10.1$ $6.35$ $4.14$ 2.65-2.65 $9.00$
+el-g3100eg $60.8$ $24.0$ 2.09-2.09 $42.4$ $80.8$ $100$ $10.6$ $7.10$ $4.19$ 2.88-2.88 $8.34$
+el-g3150eg $60.7$ $24.3$ 2.11-2.11 $42.0$ $81.0$ $150$ $11.1$ $7.15$ $4.18$ 3.05-3.05 $8.30$
 Table 2: The spectra of hadrons are in units of $\mathrm{MeV}$.
@@
 The constraints of $f_{0}(1370)$ and $a_{0}(1450)$ are combined together, due to $\rm U(3)_{V}$ symmetry.
-The definition of the parameter set is shown in Tab.
-1.
+The definition of the parameter set is shown in Table 1.
 $m_{N}$ $m_{\sigma}$ $m_{f_{0}(a_{0})}$ $m_{f_{0}^{\prime}(a_{0}^{\prime})}$ $m_{\sigma^{\prime}}$ $m_{\rho}$ $m_{\rho}$
-Empirical $938$-$940$ $400$-$800$ $960$-$1010$ $1250$-$1500$ $1430$-$1530$ $761$-$765$ $782$-$783$
+Empirical 938938-$940$ 400400-$800$ 960960-$1010$ 12501250-$1500$ 14301430-$1530$ 761761-$765$ 782782-$783$
 el-g30g $939$ $522$ $989$ $1480$ $1510$ $763$ $783$
@@
 The $n_{0}$ is in units of $\rm fm^{-3}$ and the other properties are in units of $\rm MeV$.
-The definition of the parameter set is shown in Tab.
-1.
+The definition of the parameter set is shown in Table 1.
 $n_{0}$ $E_{0}$ $E_{\mathrm{sym}}(n_{0})$ $J(n_{0})$ $L(n_{0})$ $K(n_{0})$
-Empirical $0.155\pm 0.050$ $-16.0\pm 1.0$ $31.7\pm 3.2$ $-200\pm 600$ $58.7\pm 28.1$ $250\pm 50$
-el-g30g $0.155$ $-14.6$ $30.9$ $451$ $83.6$ $418$
-el-g30e $0.155$ $-14.6$ $31.6$ $479$ $85.8$ $419$
-el-g30eg $0.155$ $-14.6$ $30.1$ $421$ $82.2$ $415$
-el-g350eg $0.155$ $-15.2$ $30.9$ $-392$ $80.7$ $370$
-el-g3100eg $0.155$ $-15.4$ $31.4$ $-1020$ $71.7$ $317$
-el-g3150eg $0.155$ $-15.6$ $31.6$ $-1470$ $63.7$ $253$
-It can be seen from Tab.
-2 that all the parameter sets reproduce hadron spectra consistent with the results from the Particle Data Group Navas et al.
+Empirical $0.155\pm 0.050$ $16.0\pm 1.0$ $31.7\pm 3.2$ $200\pm 600$ $58.7\pm 28.1$ $250\pm 50$
+el-g30g $0.155$ 14.6-14.6 $30.9$ $451$ $83.6$ $418$
+el-g30e $0.155$ 14.6-14.6 $31.6$ $479$ $85.8$ $419$
+el-g30eg $0.155$ 14.6-14.6 $30.1$ $421$ $82.2$ $415$
+el-g350eg $0.155$ 15.2-15.2 $30.9$ 392-392 $80.7$ $370$
+el-g3100eg $0.155$ 15.4-15.4 $31.4$ 1020-1020 $71.7$ $317$
+el-g3150eg $0.155$ 15.6-15.6 $31.6$ 1470-1470 $63.7$ $253$
+It can be seen from Table 2 that all the parameter sets reproduce hadron spectra consistent with the results from the Particle Data Group Navas et al.
 (2024).
-Regarding the NM properties at $n_{0}$ shown in Tab.
-3, the saturation density is consistently fixed at $0.155~{}\rm fm^{-3}$, and $E_{\rm sym}$ aligns with the empirical value of $(30.9\pm 1.9)~{}\rm MeV$.
+Regarding the NM properties at $n_{0}$ shown in Table 3, the saturation density is consistently fixed at $0.155~\rm fm^{-3}$, and $E_{\rm sym}$ aligns with the empirical value of $(30.9\pm 1.9)~\rm MeV$.
 However, the other properties decrease with the magnitude of $\tilde{g}_{3}$.
@@
 The M-R relations of NSs are shown in Fig.
-1, which is computed using the Tolman-Oppenheimer-Volkoff (TOV) equation Tolman (1939); Oppenheimer and Volkoff (1939) with pure hadron EOSs from different models.
+1, which is computed using the Tolman-Oppenheimer-Volkoff equation Tolman (1939); Oppenheimer and Volkoff (1939) with pure hadron EOS from different models.
 (a) Comparison among TM1 Sugahara and Toki (1994), FSU-$\delta 6.7$ Li et al. (2022), and el-g30eg.
@@
 1(a), the model without the $a_{0}$ meson cannot reach the GW170817 constraint due to its overly loose structure of NSs.
-The models with the $a_{0}$ meson, both el-g30eg and FSU-$\delta 6.7$, can satisfy the GW170817 and MSP J0740+6620 constraints simultaneously, highlighting the importance of the $a_{0}$ meson—an iso-vector scalar meson in NS studies.
+The models with the $a_{0}$ meson, both el-g30eg and FSU-$\delta 6.7$, can satisfy the GW170817 and MSP J0740+6620 constraints simultaneously, highlighting the importance of the $a_{0}$ meson—an isovector scalar meson in NS studies.
 From Fig.
-1(b), it is evident that only EOSs with similar contributions from the $e$ and $g$ terms can meet the observational constraint of MSP J0740+6620; otherwise, the $M_{\rm max}$ will be slightly lower.
+1(b), it is evident that only EOS with similar contributions from the $e$ and $g$ terms can meet the observational constraint of MSP J0740+6620; otherwise, the $M_{\rm max}$ will be slightly lower.
 However, the properties of NM at $n_{0}$ are not sensitive to the contributions of the $e$ and $g$ terms.
@@
 ${}+g_{\rho NN}\bar{N}\rho\tau_{3}\gamma_{0}N,$
-where $N=\binom{p}{n}$ and $\tau_{3}$ is the 3rd component of Pauli matrix. The values of the couplings are list in Table 4.
+where $N=\binom{p}{n}$ and $\tau_{3}$ is the third component of Pauli matrix. The values of the couplings are list in Table 4.
 Table 4: The values of meson-baryon coupling in various of models.
-$g_{\sigma NN}$ $g_{aNN}$ $g_{fNN}$ $g_{\omega NN}$ $g_{\rho NN}$
-TM1 Sugahara and Toki (1994) $-10.0$ $-12.6$ $-4.63$
-FSU-$\delta$6.7 Li et al. (2022) $10.2$ $6.7$ $-13.4$ $-7.27$
-el-g30g $-6.17$ $5.13$ $2.95$ $-6.09$ $5.30$
-el-g30e $-6.20$ $-5.03$ $3.00$ $6.09$ $5.30$
-el-g30eg $-5.97$ $-0.671$ $2.68$ $6.06$ $3.45$
-el-g350eg $-6.12$ $-0.852$ $2.85$ $6.37$ $3.71$
-el-g3100eg $-6.36$ $-0.442$ $3.19$ $6.73$ $3.85$
-el-g3150eg $-6.38$ $-0.413$ $3.20$ $7.09$ $4.04$
+$g_{\sigma NN}$ $g_{{a_{0}NN}}$ $g_{fNN}$ $g_{\omega NN}$ $g_{\rho NN}$
+TM1 Sugahara and Toki (1994) 10.0-10.0 12.6-12.6 4.63-4.63
+FSU-$\delta$6.7 Li et al. (2022) $10.2$ $6.7$ 13.4-13.4 7.27-7.27
+el-g30g 6.17-6.17 $5.13$ $2.95$ 6.09-6.09 $5.30$
+el-g30e 6.20-6.20 5.03-5.03 $3.00$ $6.09$ $5.30$
+el-g30eg 5.97-5.97 0.671-0.671 $2.68$ $6.06$ $3.45$
+el-g350eg 6.12-6.12 0.852-0.852 $2.85$ $6.37$ $3.71$
+el-g3100eg 6.36-6.36 0.442-0.442 $3.19$ $6.73$ $3.85$
+el-g3150eg 6.38-6.38 0.413-0.413 $3.20$ $7.09$ $4.04$
 In our bELSM, the magnitudes of the $e$ and $g$ terms determine $g_{aNN}$, while $g_{\sigma NN}$ and $g_{fNN}$ remain almost the same across different parameter sets.
@@
 This is not the case in FSU-$\delta 6.7$, which does not consider $f_{0}$.
-In bELSM, all possible terms involving the $\delta$ meson at the lowest order are considered, including all three- and four-point interactions: self-couplings, couplings to vector mesons, and couplings to scalar mesons.
+In bELSM, all possible terms involving the $a_{0}$} meson at the lowest-order are considered, including all three- and four-point interactions: self-couplings, couplings to vector mesons, and couplings to scalar mesons.
 These terms become increasingly important with density, especially those relate to vector mesons.
-By including all possible multi-meson interactions at the lowest order, respecting the chiral symmetry pattern at low energies, the $g_{\sigma NN}$ and $g_{\omega NN}$ at densities are slightly smaller than those in Walecka-type models Holinde and Machleidt (1975); Erkelenz (1974); Nagels et al.
+By including all possible multimeson interactions at the lowest-order, respecting the chiral symmetry pattern at low energies, the $g_{\sigma NN}$ and $g_{\omega NN}$ at densities are slightly smaller than those in Walecka-type models Holinde and Machleidt (1975); Erkelenz (1974); Nagels et al.
 (1978); Machleidt (1989); Sugahara and Toki (1994); Li et al.
@@
 (2021).
-However, the $E_{\rm sym}(n)$ from TM1 is almost linearly dependent on the density, which is too stiff at intermediate densities, leading to a too loose NS compared to the observations of GW170817, as shown in Fig.
+However, the $E_{\rm sym}(n)$ from TM1 is almost linearly dependent on the density, which is too stiff at intermediate densities, leading to too-loose NS compared to the observations of GW170817, as shown in Fig.
 1(a).
-This highlights the necessity of including the $a_{0}$ meson in the NS EOS, resulting in a plateau-like structure of $E_{\rm sym}(n)$ at intermediate densities.
+This highlights the necessity of including the $a_{0}$ meson in the NS EOS, resulting in a plateaulike structure of $E_{\rm sym}(n)$ at intermediate densities.
 However, a large $g_{aNN}$ will lead to a decreasing $E_{\rm sym}(n)$ at high densities, as shown in Fig.
@@
 Figure 3: The $v_{s}^{2}$ from different cases. The conformal limit is due to the asymptotic freedom of QCD at high energies.
-It can be observed that a large $\tilde{g}_{3}$ makes $v_{s}^{2}$ softer compared to $g_{aNN}$ suppression, as seen in Fig.
-3(b) and Fig.
-3(c).
+It can be observed that a large $\tilde{g}_{3}$ makes $v_{s}^{2}$ softer compared to $g_{{a_{0}NN}}$ suppression, as seen in Figs.
+3(b) and 3(c).
 The soft SV behavior, especially when not approaching the conformal limit throughout the densities, predicts a small $M_{\rm max}$ for NSs, which cannot meet the constraints of MSP J0740+6620.
-Now, let’s examine the results in Fig. 1(a):
+Now, let us examine the results in Fig. 1(a):
   • The SV of TM1 is the stiffest throughout the densities, but the predicted NS is the loosest.
@@
 Moreover, considering the SV behavior at high densities, el-g30eg allows the SV to approach the conformal limit.
-This technically stems from including all possible terms about the related baryons and mesons at the lowest order, but the deeper reason may be its close relation to QCD.
+This technically stems from including all possible terms about the related baryons and mesons at the lowest-order, but the deeper reason may be its close relation to QCD.
 All the M-R relations in this work are calculated with parameters determined by the NM properties at $n_{0}$, but the parameter space and the resulting NS structures can be further improved via statistical analysis, including constraints from NS observations, as discussed in our past work Guo et al.
@@
 II.3 Constituents of the scalar mesons
-The above discussions suggest possible magnitudes of the 2-quark and 4-quark constituents of the scalar mesons.
+The above discussions suggest possible magnitudes of the two-quark and four-quark constituents of the scalar mesons.
 We list the percentage of the two-quark component $p_{2}=\left|\cos^{2}\theta\right|$ and four-quark component $p_{4}=\left|\sin^{2}\theta\right|$ defined by Eq.
-(8) in Tab.
-5.
-Table 5: The possible percentage of the 2-quark and 4-quark constituents of the scalar mesons.
-$\sigma$ $\delta$
+(8) in Table 5.
+Table 5: The possible percentage of the two-quark and four-quark constituents of the scalar mesons.
+$\sigma$ $a_{0}$}
 $p_{4}$ $p_{2}$ $p_{4}$ $p_{2}$
@@
 el-g30e $51.2\%$ $48.7\%$ $77.2\%$ $22.8\%$
-el-g30eg $54.2\%$ $45.8\%$ $81.5\%$ $18.5\%$
+el-g30eg $54.2\%$ $45.8\%$ $\% 18.518.5\%$
 el-g350eg $52.4\%$ $47.6\%$ $77.3\%$ $20.7\%$
@@
 el-g3150eg $48.9\%$ $51.1\%$ $74.3\%$ $25.7\%$
-It is found that the $\sigma$ meson, a flavor singlet, is composed of 2-quark and 4-quark constituents in similar proportions, while the $\delta$ meson, a flavor octet, is primarily composed of 4-quark constituents.
-This conclusion is drawn after determining the bare masses of the relevant hadrons in Tab.
-2 and the NM properties in Tab.
-3.
+It is found that the $\sigma$ meson, a flavor singlet, is composed of two-quark and four-quark constituents in similar proportions, while the $a_{0}$} meson, a flavor octet, is primarily composed of four-quark constituents.
+This conclusion is drawn after determining the bare masses of the relevant hadrons in Table 2 and the NM properties in Table 3.
 This finding is consistent with previous analyses in vacuum Fariborz et al.
 (2008).
-When increasing the magnitude of the $\tilde{g}_{3}$ term, the 4-quark constituent percentages of both the $\sigma$ and $\delta$ mesons decrease to find an optimal parameter set that meets empirical constraints.
-Conversely, when the contributions of the $e$ and $g$ terms are made similar, the 4-quark constituent percentages of both the $\sigma$ and $\delta$ mesons increase.
-Regarding the parameter space found in our previous work Ma and Ma (2024), the current optimal parameter spaces are more favored, not only for their consistent NM properties at $n_{0}$ and reasonable EOS behaviors over a wider density region, as discussed above, but also for their prediction of the pion decay constant $f_{\pi}$.
-The pion decay constant can be obtained using the PCAC relation in the chiral limit (no explicit symmetry breaking due to quark mass) Fariborz et al.
+When increasing the magnitude of the $\tilde{g}_{3}$ term, the four-quark constituent percentages of both the $\sigma$ and $a_{0}$} mesons decrease to find an optimal parameter set that meets empirical constraints.
+Conversely, when the contributions of the $e$ and $g$ terms are made similar, the four-quark constituent percentages of both the $\sigma$ and $a_{0}$} mesons increase.
+Regarding the parameter space found in our previous work Ma and Ma (2024), the current optimal parameter spaces are more favored, not only for their consistent NM properties at $n_{0}$ and reasonable EOS behaviors over a wider density region, as discussed above but also for their prediction of the pion decay constant $f_{\pi}$.
+The pion decay constant can be obtained using the partially conserved axial current (PCAC) relation in the chiral limit (no explicit symmetry breaking due to quark mass) Fariborz et al.
 (2008): $f_{\pi}=2\sqrt{\alpha^{2}+\beta^{2}}$.
@@
 Effectively, we define the coupling through
-$\mathcal{L}_{\rm RMF}$ $=$ $C_{\sigma\delta^{2}}\sigma a_{0}^{2}+C_{\sigma^{2}\delta^{2}}\sigma^{2}a_{0}^{2}.$ (10)
-Then, according to Tab. 1 and Lagrangian (3b), the corresponding values of the couplings are listed in Tab. 6 for comparison.
-Table 6: The values of $C_{\sigma\delta^{2}}$ and $C_{\sigma^{2}\delta^{2}}$ in different models.
-$C_{\sigma\delta^{2}}$(MeV) $C_{\sigma^{2}\delta^{2}}$
-Zabari-19 Zabari et al. (2019) ±$\pm 1.77$$\pm 0.004$
+$\mathcal{L}_{\rm RMF}$ $=$ $C_{\sigma{a_{0}}^{2}}\sigma a_{0}^{2}+C_{\sigma^{2}{a_{0}}^{2}}\sigma^{2}a_{0}^{2}.$ (10)
+Then, according to Table 1 and Lagrangian (3b), the corresponding values of the couplings are listed in Table 6 for comparison.
+Table 6: The values of $C_{\sigma{a_{0}}^{2}}$ and $C_{\sigma^{2}{a_{0}}^{2}}$ in different models.
+$C_{\sigma{a_{0}}^{2}}$(MeV) $C_{\sigma^{2}{a_{0}}^{2}}$
+Zabari-19 Zabari et al. (2019) $\pm 1.77$ $\pm 0.004$
 FSU-$\delta$6.7 Li et al. (2022) $2.63$
-el-g30g $-1860$ $-9.40$
-el-g30e $-1940$ $-9.71$
-el-g30eg $-1480$ $-7.70$
-el-g350eg $-1690$ $-8.75$
-el-g3100eg $-2190$ $-11.0$
-el-g3150eg $-2160$ $-11.0$
+el-g30g 1860-1860 9.40-9.40
+el-g30e 1940-1940 9.71-9.71
+el-g30eg 1480-1480 7.70-7.70
+el-g350eg 1690-1690 8.75-8.75
+el-g3100eg 2190-2190 11.0-11.0
+el-g3150eg 2160-2160 11.0-11.0
 It can be seen that the magnitude of couplings between $a_{0}$ and $\sigma$ mesons are all large, differing significantly from the results of FSU-$\delta 6.7$ and Zabari-19, which are used for dense NM studies.
-The larger magnitude of the couplings in Tab.
-6 indicates that these multi-meson interactions contribute more to the EOS at both saturation density and high densities, and links the behavior of both regions more closely.
+The larger magnitude of the couplings in Table 6 indicates that these multimeson interactions contribute more to the EOS at both saturation density and high densities, and links the behavior of both regions more closely.
 This is the reason why we must increase the $K(n_{0})$ to achieve a reasonable NS structures, as shown Fig.
 1(c).
-While the parameter spaces of FSU-$\delta 6.7$ and Zabari-19 make NM properties at $n_{0}$ do not change obviously after introducing these scalar couplings, they affect the EOS at high densities much.
+While the parameters from FSU-$\delta$6.7 and Zabari-19 make NM properties at $n_{0}$ do not change obviously compared to those without couplings between the $a_{0}$ and $\sigma$ mesons, for the small magnitudes of these couplings, they affect the EOS at high densities much.
 This leads to the fact that they can handle the NM properties at $n_{0}$ and NS structures separately, compared to the bELSM.
-However, since the scalar meson couplings in $\mathcal{L}_{\rm M}$ are determined by $\alpha$, $\beta$, $e_{3}$, and $c_{4}$, which are closely related to the quark constituents of scalar mesons listed in Tab.
-5 and important physical values, such as $f_{\pi}$.
+However, since the scalar meson couplings in $\mathcal{L}_{\rm M}$ are determined by $\alpha$, $\beta$, $e_{3}$, and $c_{4}$, which are closely related to the quark constituents of scalar mesons listed in Table 5 and important physical values, such as $f_{\pi}$.
 In contrast, the Walecka-type models, as we are addressing, only introduce possible freedoms and choose operators artificially, which cannot be directly linked to QCD.
-Therefore, the results in Tab.
-6 highlight the importance of considering the chiral symmetry pattern of QCD at low energies in phenomenological analysis and including operators systematically, especially for phenomenological studies at high densities.
-III Conclusion and discussion
+Therefore, the results in Table 6 highlight the importance of considering the chiral symmetry pattern of QCD at low energies in phenomenological analysis and including operators systematically, especially for phenomenological studies at high densities.
+III Conclusion
 In this work, we continued to study NM properties and NS structures by extending the bELSM previousely proposed in Ref.
 Ma and Ma (2024) to include a new type of flavor structure which is constructed with an antisymmetric tensor of the corresponding baryon and meson representations.
-With this extension, the bELSM at the lowest order could reproduce the empirical spectra of the scalar mesons and nucleons, as well as the NM properties at saturation density.
-Explicitly, we found that without introducing the 12-quark terms as in the previous work, the reasonable behavior of nuclear matter at high densities can be well reproduced.
+With this extension, the bELSM at the lowest-order could reproduce the empirical spectra of the scalar mesons and nucleons as well as the NM properties at saturation density.
+Explicitly, we found that without introducing the two-quark terms as in the previous work, the reasonable behavior of nuclear matter at high densities can be well reproduced.
 The couplings related to $\rho$ and $\omega$ mesons can be separated due to the new flavor structures, allowing the $E_{\rm sym}$ plateau to appear at the expected density range, around $2n_{0}$, instead of below $n_{0}$ as in the previous work.
@@
 Several parameter sets were found to meet the NM properties at $n_{0}$.
-The differences lie in the magnitude of $g_{aNN}$ and the four-vector meson coupling, $\tilde{g}_{3}$.
-It was found that small $g_{aNN}$ and $\tilde{g}_{3}$ lead to more compact NSs, consistent with the constraints of MSP J0740+6620 and GW170817.
-The influences of the EOS behaviors on NS M-R relations were also discussed: a soft $E_{\rm sym}$ at intermediate densities, with a stiff behavior at subsaturation densities, is necessary for the compactness of NSs to meet the GW170817 constraints; a stiff SV behavior at high densities is necessary for the $M_{\rm max}$ of NSs to meet the MSP J0740+6620 constraints.
-Besides the well-produced NS structures from el-g30eg and FSU-$\delta 6.7$, the SV behavior of el-g30eg is more favored, as its SV behavior can approach the conformal limit at high densities, expected for the fact that QCD is asymptoticly free at high energies.
-The percentages of 2-quark and 4-quark constituents making up the scalar mesons were also calculated.
-The $\sigma$ meson was found to be composed of 2-quark and 4-quark constituents in similar proportions, while the $a_{0}$ meson is mainly composed of 4-quark constituents in the current optimal parameter spaces.
+The differences lie in the magnitude of $g_{{a_{0}NN}}$ and the four-vector meson coupling, $\tilde{g}_{3}$.
+It was found that small $g_{{a_{0}NN}}$ and $\tilde{g}_{3}$ lead to more compact NSs, consistent with the constraints of MSP J0740+6620 and GW170817.
+The influences of the EOS behaviors on NS M-R relations were also discussed: a soft $E_{\rm sym}$ at intermediate densities, with a stiff behavior at subsaturation densities, is necessary for the compactness of NSs to meet the GW170817 constraints, and a stiff SV behavior at high densities is necessary for the $M_{\rm max}$ of NSs to meet the MSP J0740+6620 constraints.
+Besides the well-produced NS structures from el-g30eg and FSU-$\delta 6.7$, the SV behavior of el-g30eg is more favored, as its SV behavior can approach the conformal limit at high densities, expected for the fact that QCD is asymptotically free at high energies.
+The percentages of two-quark and four-quark constituents making up the scalar mesons were also calculated.
+The $\sigma$ meson was found to be composed of two-quark and four-quark constituents in similar proportions, while the $a_{0}$ meson is mainly composed of four-quark constituents in the current optimal parameter spaces.
 This is consistent with previous analyses in vacuum Fariborz et al.
 (2008).
-The variation of the constituents due to different magnitudes of $g_{aNN}$ and four-vector meson couplings were also provided for future studies on scalar mesons and the macroscopic properties of compact objects in the universe.
-However, the quark constituents make the couplings related to $a_{0}$, both types: $g_{a_{0}NN}$ and $C_{\sigma^{i}\delta^{2}}$, different from the results of FSU-$\delta 6.7$ and Zabari-19, which are used for the symmetry energy at high densities.
+The variation of the constituents due to different magnitudes of $g_{{a_{0}NN}}$ and four-vector meson couplings were also provided for future studies on scalar mesons and the macroscopic properties of compact objects in the universe.
+However, the quark constituents make the couplings related to $a_{0}$, both types: $g_{{a_{0}NN}}$ and $C_{\sigma^{i}{a_{0}}^{2}}$, with i=1, 2i=1,\ 2, different from the results of FSU-$\delta 6.7$ and Zabari-19, which are used for the symmetry energy at high densities.
 This again signifies the importance of considering the chiral symmetry pattern of QCD at low energies in phenomenological analyses beyond the saturation density.
@@
 (2024a, b) focus on the relationships between the QCD symmetry pattern and compact objects in the universe.
-In the near future, improved AI analysis on a general Walecka-type model with the inclusion of all types of lowest-lying mesons and baryons will be available, which will help further investigate bELSM at densities due to its numerical complexity.
+In the near future, improved AI analysis on a general Walecka-type model with the inclusion of all types of lowest-lying mesons and baryons will be available, which will help further investigate bELSM at finite densities due to its numerical complexity.
 Additionally, studies based on bELSM with the inclusion of strange freedoms are underway to address the properties of compact objects with hyperons.
-Meanwhile, the Hartree-Fock approach with $\chi\rm EFT$ will also be initially completed to study the EOS of NM, including quantum effects, which are ignored in the current RMF approximation.
+Meanwhile, the Hartree-Fock approach with chiral effective field theories will also be initially completed to study the EOS of NM, including quantum effects, which are ignored in the current RMF approximation.
 Acknowledgements.
@@
 ZXL2024363.
-Appendix A The lowest order bELSM
-In this appendix, we provide details of the bELSM at the lowest order before the RMF approximation.
+Appendix A The lowest-order bELSM
+In this Appendix, we provide details of the bELSM at the lowest-order before the RMF approximation.
 In the model, the scalar mesons $S$ and pseudoscalar mesons $P$ are combined into a compact form $\Phi=S+iP$:
 $\Phi=S+iP=\left(\begin{array}[]{ccc}\frac{\left(\sigma_{N}+a_{0}^{0}\right)+i\left(\eta_{N}+\pi^{0}\right)}{\sqrt{2}}&a_{0}^{+}+i\pi^{+}&K_{S}^{+}+iK^{+}\\ a_{0}^{-}+i\pi^{-}&\frac{\left(\sigma_{N}-a_{0}^{0}\right)+i\left(\eta_{N}-\pi^{0}\right)}{\sqrt{2}}&K_{S}^{0}+iK^{0}\\ K_{S}^{-}+iK^{-}&\bar{K}_{S}^{0}+i\bar{K}^{0}&\sigma_{S}+i\eta_{S}\end{array}\right)\ .$ (11)
-The representations of the 2-quark and 4-quark (pseudo-)scalar configurations are similar, but with ”prime” and ”hat-prime” indices, as defined in the main text.
-For the vector and axial-vector mesons, they are expressed as:
+The representations of the two-quark and four-quark (pseudo-)scalar configurations are similar, but with ”prime” and ”hat-prime” indices, as defined in the main text.
+For the vector and axial-vector mesons, they are expressed as,
 $(R,L)^{\mu}=V^{\mu}\pm A^{\mu}=\frac{1}{\sqrt{2}}\left(\begin{array}[]{ccc}\frac{\omega_{N}^{\mu}+\rho^{\mu 0}}{\sqrt{2}}\pm\frac{f_{1N}^{\mu}+a_{1}^{\mu 0}}{\sqrt{2}}&\rho^{\mu+}\pm a_{1}^{\mu+}&K^{*\mu+}\pm K_{1}^{\mu+}\\ \rho^{\mu-}\pm a_{1}^{\mu-}&\frac{\omega_{N}^{\mu}-\rho^{\mu 0}}{\sqrt{2}}\pm\frac{f_{1N}^{\mu}-a_{1}^{\mu 0}}{}&K^{*\mu 0}\pm K_{1}^{\mu 0}\\ K^{*\mu-}\pm K_{1}^{\mu-}&\bar{K}^{*\mu 0}\pm\bar{K}_{1}^{\mu 0}&\omega_{S}^{\mu}\pm f_{1S}^{\mu}\end{array}\right)\ ,$ (12)
-where $L(R)^{\mu}$ stands for the left(right)-handed vector mesons, and $V(A)^{\mu}$ denotes the vector(axial-vector) mesons. The baryon octet has expression:
+where $L(R)^{\mu}$ stands for the left-(right)-handed vector mesons, and $V(A)^{\mu}$ denotes the vector(axial-vector) mesons. The baryon octet has expression:
 $B=\left(\begin{array}[]{ccc}\frac{\Lambda}{\sqrt{6}}+\frac{\Sigma^{0}}{\sqrt{2}}&\Sigma^{+}&p\\ \Sigma^{-}&\frac{\Lambda}{\sqrt{6}}-\frac{\Sigma^{0}}{\sqrt{2}}&n\\ \Xi^{-}&\Xi^{0}&-\frac{2\Lambda}{\sqrt{6}}\end{array}\right)\ .$ (13)
-The chiral symmetry transformations for (pseudo-)scalar configurations are given by Fariborz et al. (2005, 2008, 2009):
+The chiral symmetry transformations for (pseudo)scalar configurations are given by Fariborz et al. (2005, 2008, 2009):
 $\Phi^{\prime}\rightarrow g_{\mathrm{L}}\Phi^{\prime}g_{\mathrm{R}}^{\dagger},\quad\hat{\Phi}^{\prime}\rightarrow g_{\mathrm{L}}\hat{\Phi}^{\prime}g_{\mathrm{R}}^{\dagger}\ ,$ (14)
-where $g_{\mathrm{L},\mathrm{R}}\in\mathrm{SU}(3)_{\mathrm{L},\mathrm{R}}$. They remain unchanged for $\rm U(1)_{V}$ transformation. But for $\rm U(1)_{A}$ transformation, they transform as:
+where $g_{\mathrm{L},\mathrm{R}}\in\mathrm{SU}(3)_{\mathrm{L},\mathrm{R}}$. They remain unchanged for $\rm U(1)_{V}$ transformation. But for $\rm U(1)_{A}$ transformation, they transform as,
 $\Phi^{\prime}\rightarrow e^{2i\nu}\Phi^{\prime},\quad\hat{\Phi}^{\prime}\rightarrow e^{-4i\nu}\hat{\Phi}^{\prime}\ ,$ (15)
-with $\nu$ being the chiral angle, and $q\bar{q}$ and $(qq)(\bar{q}\bar{q})$ configurations are assumed for 2-quark and 4-quark components, respectively.
+with $\nu$ being the chiral angle, and $q\bar{q}$ and $(qq)(\bar{q}\bar{q})$ configurations are assumed for two-quark and four-quark components, respectively.
 For the vector and axial-vector mesons, their chiral transformations are Parganlija et al.
-(2013):
+(2013),
 $L_{\mu}\rightarrow g_{\mathrm{L}}L_{\mu}g_{\mathrm{L}}^{\dagger},\quad R_{\mu}\rightarrow g_{\mathrm{R}}R_{\mu}g_{\mathrm{R}}^{\dagger}\ ,$ (16)
@@
 The baryons are denoted as $N_{\mathrm{R}}^{(\mathrm{RR})}$, $N_{\mathrm{L}}^{(\mathrm{LL})}$, $N_{\mathrm{L}}^{(\mathrm{RR})}$, and $N_{\mathrm{R}}^{(\mathrm{LL})}$, where the superscripts represent the quark configuration and the subscripts represent the diquark configuration.
-The chiral transformations for baryons are:
+The chiral transformations for baryons are
 $N_{\mathrm{R}}^{(\mathrm{RR})}$ $\rightarrow$ $g_{\mathrm{R}}N_{\mathrm{R}}^{(\mathrm{RR})}g_{\mathrm{R}}^{\dagger}\ ,\quad N_{\mathrm{L}}^{(\mathrm{RR})}\rightarrow g_{\mathrm{L}}N_{\mathrm{L}}^{(\mathrm{RR})}g_{\mathrm{R}}^{\dagger}\ ,$
 $N_{\mathrm{R}}^{(\mathrm{LL})}$ $\rightarrow$ $g_{\mathrm{R}}N_{\mathrm{R}}^{(\mathrm{LL})}g_{\mathrm{L}}^{\dagger}\ ,\quad N_{\mathrm{L}}^{(\mathrm{LL})}\rightarrow g_{\mathrm{L}}N_{\mathrm{L}}^{(\mathrm{LL})}g_{\mathrm{L}}^{\dagger}\ .$ (17)
-And the $\rm U(1)_{A}$ transformations are:
+And the $\rm U(1)_{A}$ transformations are
 $N_{\mathrm{R}}^{(\mathrm{RR})}$ $\rightarrow$ $e^{-3iv}N_{\mathrm{R}}^{(\mathrm{RR})}\ ,\quad N_{\mathrm{L}}^{(\mathrm{RR})}\rightarrow e^{-iv}N_{\mathrm{L}}^{(\mathrm{RR})}\ ,$
@@
 with excited state $N^{*}$ being neglected since it is irrelevant to the present work.
-Before constructing the full Lagrangian, a series of power-counting rules are needed to organize the effective operators:
+Before constructing the full Lagrangian, a series of power counting rules is needed to organize the effective operators:
   • The operators are limited to dimension-4, as higher-dimensional operators are suppressed by the cutoff scale.
@@
   • The explicit symmetry breaking caused by quark mass is treated as a perturbation and is ignored in the current work.
-Consequently, the full Lagrangian at the lowest order is
+Consequently, the full Lagrangian at the lowest-order is
 $\mathcal{L}_{\rm LO}=\mathcal{L}_{\rm M}+\mathcal{L}_{\rm V}+\mathcal{L}_{\rm B}.$ (20)
-The (pseudo-)scalar meson part is Fariborz et al. (2008):
+The (pseudo)scalar meson part is Fariborz et al. (2008)
 $\mathcal{L}_{\rm M}$ $=$ $\frac{1}{2}\operatorname{Tr}\left(\partial_{\mu}\Phi^{\prime}\partial^{\mu}\Phi^{\prime\dagger}\right)+\frac{1}{2}\operatorname{Tr}\left(\partial_{\mu}\hat{\Phi}^{\prime}\partial^{\mu}\hat{\Phi}^{\prime\dagger}\right)+c_{2}\operatorname{Tr}\left(\Phi^{\prime}\Phi^{\prime\dagger}\right)-c_{4}\operatorname{Tr}\left(\Phi^{\prime}\Phi^{\prime\dagger}\Phi^{\prime}\Phi^{\prime\dagger}\right)-d_{2}\operatorname{Tr}\left(\hat{\Phi}^{\prime}\hat{\Phi}^{\prime\dagger}\right)$ (21)
@@
 And after substituting Eq. (19) into the Lagrangian, the baryon part can be rewritten as
-$\mathcal{L}_{\rm B}$ $=$ ${\rm Tr}\left(\bar{B}i\not{\partial}B\right)+c{\rm Tr}\left(\bar{B}\gamma_{\mu}V^{\mu}B+\bar{B}\gamma_{\mu}\gamma_{5}A^{\mu}B\right)+c^{\prime}{\rm Tr}\left(\bar{B}\gamma_{\mu}BV^{\mu}+\bar{B}\gamma_{\mu}\gamma_{5}BA^{\mu}\right)$ (24)
+$\mathcal{L}_{\rm B}$ $=$ ${\rm Tr}\left(\bar{B}i\not{\partial}B\right)+c{\rm Tr}\left(\bar{B}\gamma_{\mu}V^{\mu}B+\bar{B}\gamma_{\mu}\gamma_{5}A^{\mu}B\right)+{c^{\prime}{\rm Tr}\left(\bar{B}\gamma_{\mu}BV^{\mu}\right)+c^{\prime}_{A}\left(\bar{B}\gamma_{\mu}\gamma_{5}BA^{\mu}\right)}$ (24)
 ${}+h\epsilon_{abc}\epsilon^{def}\left[\left(\bar{B}\right)_{ad}\gamma_{\mu}\left(B\right)_{be}\left(V^{\mu}\right)_{cf}+\left(\bar{B}\right)_{ad}\gamma_{\mu}\gamma_{5}\left(B\right)_{be}\left(A^{\mu}\right)_{cf}\right]$
@@
 ${}-\frac{e}{2}\epsilon_{abc}\epsilon^{def}\left[\left(\bar{B}\right)_{ad}\left(\Phi^{\prime}+\Phi^{\prime\dagger}\right)_{be}\left(B\right)_{cf}+\left(\bar{B}\right)_{ad}\gamma_{5}\left(\Phi^{\prime}-\Phi^{\prime\dagger}\right)_{be}\left(B\right)_{cf}\right]\ ,$
-where $c=(c_{1}+c_{2})/2$, $c^{\prime}=(c_{3}+c_{4})/2$, $h=n_{a}/2$, $g=-n_{g}$ and $e=-n_{e}$.
+where $c=(n_{1}+n_{2})/2$, $c^{\prime}=(n_{3}+n_{4})/2$, $c^{\prime}_{A}=(n_{3}-n_{4})/2$, $h=n_{a}/2$, $g=-n_{g}$ and $e=-n_{e}$.
 Appendix B Approximations in the current work
@@
 Additionally, the system composed of these baryons is considered to be infinite, homogeneous, and static, so that any quantum numbers and currents are averaged to zero.
-This results in the following approximations
+This results in the approximations
 $\langle\omega^{\mu}\rangle$ $=$ $\langle\omega^{0}\rangle\delta^{\mu,0}=\omega\delta^{\mu,0}\ ,$
@@
 $\langle\pi^{i}\rangle$ $=$ $0\ ,$ (25)
-and other related meson fields are also treated similarly. For instance, $K$ mesons are excluded due to their non-zero strange quantum number.
-Mesons heavier than $1\rm GeV$, such as $a^{\prime}_{0}$ and $\omega_{S}$, are ignored since their contributions are suppressed at the current working scale.
+and other related meson fields are also treated similarly. For instance, $K$ mesons are excluded due to their nonzero strange quantum number.
+Mesons heavier than $1~\rm GeV$, such as $a^{\prime}_{0}$ and $\omega_{S}$, are ignored since their contributions are suppressed at the current working scale.
 This leads to the following parameter choices in the current work: $\tilde{g}_{3}=2(g_{3}+g_{4})$, $\tilde{h}_{2}=2h_{1}+h_{2}$, while the other parameters remain the same as the original ones after the approximations.
-Details on how to solve the RMF equations of motion (EOMs) to obtain the equation of state (EOS) after these approximations are provided in our other work Guo et al.
+Details on how to solve the RMF equations of motion to obtain the equation of state after these approximations are provided in our other work Guo et al.
 (2024b).